### v1

# Can Robots “Taste” Grapes? Estimating SSC with Simple RGB Sensors

Abstract
In table grape cultivation, harvesting depends on accurately assessing fruit quality. While some characteristics, like color, are visible, others, such as Soluble Solid Content (SSC), or sugar content measured in degrees Brix (°Brix), require specific tools. SSC is a key quality factor that correlates with ripeness, but lacks a direct causal relationship with color. Hyperspectral cameras can estimate SSC with high accuracy under controlled laboratory conditions, but their practicality in field environments is limited. This study investigates the potential of simple RGB sensors under uncontrolled lighting to estimate SSC and color, enabling cost-effective, robot-assisted harvesting. Over the 2021 and 2022 summer seasons, we collected grape images with corresponding SSC and color labels to evaluate algorithmic solutions for SSC estimation on embedded devices commonly used in robotics and smartphones. Our results demonstrate that SSC can be estimated from visual appearance with human-like performance. We propose computationally efficient histogram-based methods for resource-constrained robots and deep learning approaches for more complex applications.
keywords:
Precision Agriculture , SSC Estimation , Computer Vision , Robotics , RGB cameras
[…]
[inst1]organization=Sapienza University of Rome Department of Computer, Control and Management Engineering ,city=Rome, postcode=00185, state=RM, country=Italy
1 Introduction
The use of passive imaging sensors to measure agronomic properties of crops enables fast, non destructive monitoring that, when integrated with machinery or robots, can be efficiently scaled to cover entire fields. RGB cameras are commonly employed to monitor external conditions, such as detecting pests and diseases, while multi-spectral cameras are occasionally used to derive indices related to vegetation health. However, assessing internal crop properties presents a more complex challenge. Hyperspectral imagery (HSI) show potential in this area, as their ability to capture specific wavelength responses can be correlated with the absorption patterns of certain chemical properties. Nevertheless, hyperspectral cameras are typically slow to acquire images, highly sensitive to illumination conditions, and usually unsuitable for use on operational agricultural machinery.
Figure 1: Computer Vision plays a crucial role in enabling automatic harvesting operations; however, the requirements for high-quality performance often conflict with the economic and practical constraints of robot design. This work demonstrates the development and deployment of a simple RGB-based algorithm for SSC and color quality estimation, addressing the complexities of field operations while utilizing cost-effective sensors.
These competing aspects collide with the desirable tension to bring more robots on the farming fields, in order to both face the shortage of workforce, but also to manage the crops in a more precise way. In fact, many cultivations are labor-intensive, and for various economic and social reasons, it is increasingly difficult to find human workers to perform operations such as harvesting. Table grape cultivation is an example of such labor-intensive cultivation that is increasingly popular in the markets but faces a workforce shortage. This shortage would be mitigated by the introduction of robot workers that could, to some extent, assist human workers in the field. To be able to do so, the robotic platform should be able to assess the quality and ripeness of the fruit using mainly passive sensors (active manipulation for perception would be orders of magnitude slower). While solutions to equip robots with hyperspectral cameras are becoming more prevalent, their high cost and operational complexity remain major barriers to widespread adoption. So, the question of designing a passive quality estimation system using low-cost, reliable, RGB cameras, or similar devices is still open.
This work addresses this question by validating its feasibility and proposing high-performance algorithmic solutions tailored for practical deployment in agricultural settings. In the following paragraphs, we provide a brief review of the relevant literature, highlighting key contributions to contextualize this study. For a broader overview of fruit ripeness estimation methods—particularly for grapes using hyperspectral and RGB imaging—we refer the reader to [11, 17].
1.0.1 Hyperspectral applications to agricultural robotics
Robots have been used for the monitoring and quality estimation of different crops and are now finding some industrial applications in high-value crops (wine grapes [20, 12], strawberries [23, 22], blueberries [19], among others). Some authors have explored specifically the quality estimation of fruits with hyperspectral sensors. In [6] an HSI device is mounted on an AGV to measure the TSS of Tempranillo wine grapes. The AGV moves along the vineyard row acquiring spectral data in a push-broom fashion. To achieve color constancy the authors perform an indirect calibration for exposure, while illuminant calibration was performed every 5m with a white reference. The impact of these choices on the accuracy of the estimation is not clear, and the authors report an RMS on °Brix estimate as low as $1.274\textdegree$. In addition, the method requires considerable post-processing time (5 hours), making it not well suited for robotic harvesting.
[18] propose a trailing cart mounting two HSI devices to monitor raspberries. However, the measurement is taken using an overarching arm from which a white reference is hung. [16] propose the Ladybug robot equipped with an HSI device built for low horticultural crops. The robot AGV is a straddle model, where the illumination constancy is granted through the use of a stroboscopic light.
In preparation for robotic use, [9] perform a very thorough experimentation using VIS-SWIR images to estimate SSC in four wine grapes varieties. The authors analyze the contribution of different wavelengths to the estimation algorithms in depth and apply various algorithms for the prediction. Across the proposed models (Partial Last Square, Random Forest, Convolutional Neural Networks, and Support Vector Machines) and four grape varieties they obtain an RMS error ranging from 1.70 to 2.20 °Brix.
In a follow-up study, to address the problem of illuminant estimation, [15] propose to learn the illuminant normalization trough the use of an auto-encoder. The solution is similar to ours, in the sense that it learns the typical illuminant distribution, but loses the absolute metric accuracy of having a white reference in the scene. The authors explore four varieties of wine grapes, declaring RMSE values that range from 1.66 to 2.20 °Brix.
While the aforementioned approaches demonstrate the notable successes of HSI in agricultural robotics, they also exhibit several key limitations, including high sensor costs, operational complexity, lengthy post-processing times, and limited adaptability to crops where overarching machinery is impractical, such as table grapes. In contrast, the approach proposed in this work achieves comparable performance without relying on HSI, operates in real-time, and is thus better suited for integration into robotic systems.
1.0.2 RGB applications to fruit quality assessment
Given the limitations of HSI illuminant and color calibration and the cost of hyperspectral devices, many authors proposed approaches based on standard RGB sensors. [1] have explored the use of color imaging combined with advanced numerical techniques, such as interval numbers (INs), and neural network regressors to estimate grape maturity indices (e.g., TSS, TA, and pH), demonstrating the feasibility of non-destructive and cost-effective methods for integration into autonomous harvesting systems. In this work, the authors declare a considerable 6.6259 MSE Baumé, which approximately corresponds to an RMSE of 4.63 °Brix.
[21] perform a very interesting analytical study on 180 samples of Kyoho grape berries using an RGB camera. The authors propose the use of some feature indexes, similar in principle to the well-known NDVI, but computed with the standard RGB channels, obtaining a very low cross-validation error of 0.78 °Brix (RMSE) on a dataset with a standard deviation of 1.37°. The study is performed in laboratory conditions, but it is nonetheless significant for our work, as it shows the effectiveness of data-driven estimations with simple RGB cameras.
[3] proposed a non-destructive estimation technique based on an RGB camera using random forest predictors trained on color histograms. The histogram was computed in CIELab color space by clustering the colors and then choosing the ones with the highest intrinsic variance. The classification task was performed on a 5-level quality scale, and no explicit SSC measurement was included in the labeling process. The experimental setup involved 200 bunches, evenly divided between Italia and Victoria varieties, shot with artificial illumination, and image features were computed after background subtraction.
This growing body of research on RGB-based estimation of internal crop characteristics highlights both the scientific and practical interest in this approach. However, to the best of our knowledge, existing methods either fall short of achieving the performance levels of HSI imaging or fail to transition effectively from controlled laboratory conditions to field environments with uncontrolled lighting.
Building on the challenges and opportunities highlighted in previous works, this study focuses on the problem of perception for quality estimation in table grape harvesting. Accurate quality estimation can enable robots to assist human workers or even surpass their capabilities during field operations. In high-quality table grape production, harvesting decisions typically follow a two-step process: an agronomist first samples grape bunches across the vineyard using a refractometer to measure SSC in °Brix. If the °Brix is sufficiently high, additional visual characteristics, such as color intensity and distribution, are evaluated. Workers are then instructed to harvest grape bunches based solely on appearance and location within the vineyard. While °Brix and color are not causally linked, they share latent ripening factors [8], allowing human experts to approximate °Brix from color to some extent. This work aims to replicate and enhance this human capability using data-driven methods and low-cost RGB sensors.
The contributions of our study are as follows:
  • 1.
    A dataset of table grape images collected across different seasons and with different sensors, labeled with both °Brix values and color categories following the aforementioned agronomic procedure, thus enabling studies on SSC estimation in realistic field conditions.
  • 2.
    Experiments and user studies to explore the extent to which °Brix can be estimated from color alone. These insights informed the design of algorithms and highlighted key factors for future research.
  • 3.
    Two different estimation algorithms have been proposed and validated:
    1. a)
      A lightweight method based on handcrafted features for constrained computational systems.
    2. b)
      A deep, fully convolutional multi-task architecture that achieves high performance with limited training data.
  • 4.
    The methods do not require complex calibration and are designed for real-time use in robotic harvesting or mobile devices.
The paper is structured as follows: Section 2 describes the dataset that was collected and the estimation methods that were designed; Section 3 defines the experimental setup, and describes and comments on the experimental validation of the problem and the algorithms. Finally, Section 4 draws conclusions and outlines a path for future work.
2 Materials and Methods
2.1 Dataset
In this Section, we describe the data collection and labeling process.
2.1.1 Experimental Field
The experimental field is located in southern Lazio, Italy. The vineyard consists of two plots measuring approximately 114 m x 51 m (0.58 ha) and 122 m x 48 m (0.58 ha). The vineyards are structured as a traditional trellis system known as Tendone, with a wide distance between each plant (3 x 3 m). All structures are traditionally covered with plastic and netting to protect grapes from rain and hail. The plantations are all older than three years, healthy, and in full production, thus representing a typical working condition for the validation of agronomic activities such as fruit harvesting or vine pruning.
Among the varieties cultivated in this field, Black Pizzutello (Pizzutello nero), a distinct variety from the Lazio region, was selected due to its importance for the regional economy and color variability.
2.1.2 Data collection and labeling
We collected static images during two summer seasons using different camera sensors. Three different devices, across multiple years, have been used to collect the images. We employed a smartphone, a Reflex, and a Realsense d435i to represent some of the possible covariate shifts that occur between the collection of the initial data on the field by a farmer or an agronomist (smartphone), a high-quality sensor (reflex) and a low-quality robotic camera (d435i). Table 1 depicts the working categories we used in this study. A more detailed description of all the data collected and the different characteristics is given in Appendix A for the interested reader.
Figure 2: Example of an image collected with a gray card for white balance correction. An agronomist performs on-field measurements by randomly sampling the SSC of berries from the top, middle, and lower parts of the bunch and then averaging the values.
Name Split # samples SSC mean SSC std SSC MAD Description
$T1_{STD}$ 112 17.19 2.43 1.99 Phone images in JPG format captured in the 2021 season
$T1_{WB}$ 112 17.19 2.43 1.99 Phone images color corrected from RAW captured in the 2021 season
$T2_{STD}$ tr 212 16.23 2.52 2.02 Phone images captured in JPG format during the 2021 season.
$T2_{STD}$ te 92 18.91 2.16 1.71 Phone images captured in JPG format during the 2022 season.
$T2_{WB}$ tr 150 17.54 2.46 2.04 Phone images captured in RAW format during the 2021 season.
$T2_{WB}$ te 92 18.91 2.16 1.71 Phone images captured in RAW format during the 2022 season.
$T2_{ALL}$ tr 300 17.54 2.46 2.04 Phone and reflex images in RAW format, 2021 season.
$T2_{ALL}$ te 92 18.91 2.16 1.71 Phone and reflex images in RAW format, 2022 season.
$T3_{STD}$ tr 352 17.67 2.75 2.33 All jpg images, across all seasons and devices.
$T3_{STD}$ te 44 15.96 1.68 1.36 The test set contains images with SSC content close to the threshold (16° Brix).
$T3_{WB}$ tr 440 18.27 1.98 1.96 All white balanced images, across all seasons and devices.
$T3_{WB}$ te 44 15.96 1.36 1.36 The test set contains images with SSC content close to the threshold (16° Brix).
Table 1: $T1_{STD}$ and $T1_{WB}$ consist respectively in jpg and raw versions of images collected with a smartphone camera during the 2021 season. $T2$ datasets are images collected with a smartphone camera, but where the training splits are from the 2021 season, and the test set from the 2022 season. In this case, there is no exact correspondence between $T2_{STD}$ and $T2_{WB}$ jpg and raw images. The $T3$ datasets serve a different purpose, as they combine images from the $T1$ and $T2$ datasets, mixing data across all seasons and devices, specifically the smartphone, Reflex (only for the WB variant) and d435i cameras. The $T3$ training set has more samples than the $T2$. At the same time, the $T3$ test set consists of images with SSC threshold quite close to the ripeness level for harvesting (16° for Pizzutello), making it significantly more challenging than the training set.
Labels: All the images were labeled both for SSC and color. For SSC, we used an Atago Hikari PAL2 digital refractometer, capable of measuring berry sugar content (degree Brix, °Brix) by contact without damaging the berry itself. This instrument has a range from 9° to 25°. Typically, depending on the grape variety, an SSC of 16°-18° is considered good enough for harvesting. Since we are trying to assess how well the estimation based on appearance would perform compared to human performance, we adopted a labeling procedure similar to the SSC sampling performed by agronomists. Figure 2 visually illustrates the process. Every berry in a grape bunch has a different SSC, and the sugar concentration in each berry is also not uniform due to gravity. For this reason, common practice is to sample the SSC three times, from the top, middle, and bottom of the bunch, and then average the results. For consistency, the refractometer was placed on the side of the berry in each measurement. For color, a categorical classification was required. Based on agronomic production use cases, we defined a color quality chart with the help of experienced agronomists. The color scale goes from 1 to 6, with whole number increments, where 1 represents the best quality and 6 the worst. Usually, values of 1 and 2 are considered high-quality produce, while 3 is still acceptable for some markets. All the other values are considered unripe. This color index accounts for color saturation and distribution across the berries and the bunch. An example of the color variability is given in Fig: 3.
Figure 3: Examples of the six color levels for color classification. A clear distinction between categories may be challenging for an untrained eye. However, our labeling validation tests described in Section 3.1.2 indicate consistent labeling despite some uncertainty in the classification of neighboring classes.
2.2 Estimation with Histogram Features
The main use-case for the estimators we design in this work is to be deployed on a robotic platform, embedded edge devices, or smartphone-like device, thus with limited memory and computational power. For this reason, the first approach is based on classic hand-crafted features, specifically on histograms. RGB channel histograms were chosen as the basic building blocks since the problem largely depends on color intensity and distribution. To capture distribution differences, we divided each bunch image into a grid of rectangular regions and computed the RGB histogram for each region. A key observation that can be made with this setup is that even with tight bounding boxes from the detector, some regions will often contain background information due to the shape of grape bunches. To address this issue, we modified the feature processing to retain only specific regions in a ”cross-shaped” selection, as illustrated in Figure 4. This selection focuses on the areas of the image where grape content is most likely to be present.
Figure 4: Each grape bunch in the scene is cropped based on the bounding box provided by the detector. The resulting image is then binned, with some bins dropped as they contain background information with high probability. This picture shows an example, but the actual bin size and dropping schemes are hyper-parameters described in Section 2. In this simple example, the n-uple defining column and row ranges described by Eq.1 would be [1,3,1,2]
To simplify hyper-parameter exploration, we defined three approaches for region selection: a ”fat” cross-shaped pattern, a ”thin” one, and one using all regions without any exclusion. More in detail, the three configurations are defined by the starting and ending cell in the column and row direction as follows:
[…]
We compared these approaches to determine the most effective feature selection strategy. In summary, the hyper-parameters that were considered while designing the features are:
  • 1.
    Image starting size (width, height)
  • 2.
    Number of cells in both the width and height directions
  • 3.
    The cross pattern used to remove the background
  • 4.
    Use of HSV instead of RGB channels
We performed a hyper-parameter search for both the task of SSC and color degree estimation using Ridge Regression. For brevity, since color estimation is an easier task to perform, in Section 3 we omit the details of the experiments for color estimation, while the SSC ones are explored in depth.
2.3 Estimation with Multi-Task Learning DNN Architecture
As mentioned before, one of the research questions we are exploring is the correlation between the color of the berries and sugar content. The correlating latent factors are many and difficult to model explicitly. Among them are climate, weather, sunlight exposure, pruning strategies, and the use of nitrogen and other chemicals. Given this complex interaction between different factors, we decided to design a neural architecture that could extract common features to be used for both estimation tasks.
Our model is shown in Figure 5. To distill a feature vector, we employed an Auto-Encoder (AE) architecture, instead of using only a feature extraction backbone alone. The reason is that we are in a data-scarcity scenario, so having the additional reconstruction loss of the auto-encoder, focusing on appearance, helps in strengthening the back-propagation signal and reduces over-fitting. Our auto-encoder is a UNet with a ResNet18 backbone. In general, UNet-like architectures alone are not suitable for reconstruction problems, since the skip connections would cause the network to simply learn an identity function. Our architecture avoids this problem, by using the bottleneck feature output as the input for the feature extraction section. From those features, a convolutional block is used to extract the features that are then fed to the task-specific heads. The two heads have the same structure, but in the case of color estimation, the outputs are logits for the multi-class problem, while for the SSC estimation are the actual estimations.
The loss for the architecture is the combination of the reconstruction loss of the AE (pixel-wise Mean Squared Error (MSE)), with the addition of a Mean Absolute Error (MAE) loss for the SSC estimate, and a Cross-Entropy loss for the color classes:
$\mathcal{L}_{MT}=\mathcal{L}_{rec}+\alpha\mathcal{L}_{color}+\beta\mathcal{L}_{brix}$ (2)
where $\alpha$ and $\beta$ are weighting factors to balance the different contributions of the three losses, whose values have been cross-validated together with other hyper-parameters.
Figure 5: The multi-task network is based on a U-Net [13] with a ResNet18 [7] backbone, designed to encode the image into a compact representation. Features extracted from the bottleneck layer are fed into a two-head estimator: one for SSC (regression) and another for color (classification). The two heads share a common ”neck,” a CNN with 512 output channels, ensuring part of the representation is shared between the tasks. Each head has a symmetrical structure, comprising three CNN blocks that progressively reduce the feature size, followed by adaptive average pooling [24] and a 1x1 convolution [10] for the final estimation.
2.3.1 Hyper-parameter Validation
Hyperparameter selection in multi-task learning requires identifying a single metric $\mathcal{S}$ that effectively summarizes performance across different models, which is particularly challenging when tasks differ significantly.
For SSC estimation, mean absolute error (MAE) on the °Brix scale provides an intuitive measure. Color estimation, however, is better assessed using Balanced Accuracy ($A_{color}$) due to the categorical nature of the target variable (although the cross-entropy loss is also applicable it is less interpretable). We combine these two metrics to evaluate the overall performance of the estimators. However, since the metrics are in different scales and units we normalize them using a factor $\eta$.
Our evaluation criteria also align with agronomic quality standards, emphasizing the importance of avoiding the harvest of unripe bunches, while allowing for some variation in quality among ripe bunches. This is achieved by introducing hinge-like penalties that apply only when estimations fall below predefined quality thresholds for SSC ($t_{brix}$) and color ($t_{color}$) estimation, ensuring that models are primarily evaluated on their ability to meet minimum quality standards. The final score, $\mathcal{S}$ integrates the normalized performance metrics and the agronomic quality criteria.
The formal definition of the score $\mathcal{S}$ is the following:
$\mathcal{S}$ $=\eta(1-A_{color})+\mathcal{L}_{brix}+\mathcal{L}_{CH}+\mathcal{L}_{BH}$ (3)
$L_{BH}$ $=\max(0,L_{brix}-t_{brix}))^{\rho}$ (4)
$L_{CH}$ $=\max(0,-\eta(A_{color}-t_{color}))^{\rho}$ (5)
where $\rho$ is the degree of the two hinge penalties $\mathcal{L}_{BH}$ (for °Brix) and $\mathcal{L}_{CH}$ (for color), set to 2 in all our experiments. $t_{brix}$ and $t_{color}$ represent the agronomic quality thresholds for °Brix and color, respectively, and $\eta$ is the normalization factor between SSC and color scores. In our experiments, $t_{brix}=2.5$Brix, which is comparable with the standard deviation of the training data, and $t_{color}=0.8$, since an error rate of $80\%$ is usually considered acceptable for human performance in this context [14]. The normalization factor is empirically set as
$\eta=\frac{t_{brix}}{t_{color}}.$ (6)
3 Experiments
In designing the experimental setup, we had two main aims. The first one was to give an initial answer on factors are relevant in estimating SSC from appearance, in order to better understand the problem for future work and data collection campaigns. The second aim is to validate the proposed features and estimator models. This section begins with the validation of the collected dataset, followed by a description of the exploratory experiments. Finally, we describe the validation of our proposed features and models, together with the ablation studies. The data and code for all the experiments can be found on the accompanying GitHub repository [^1] To be disclosed upon publication.
3.1 Exploratory User Studies
3.1.1 Human Evaluated SSC and Color Correlation
Feature description
v1, v2, v3 Upper, middle and lower °Brix measurements.
m_brix Mean °Brix computed from v1, v2, v3.
m_b_t Rounded m_brix (unused).
b_class Binary °Brix class for harvesting (Sufficient, Insufficient). Computed from m_brix with threshold $17$°.
c_val Expert given score for color development after varaison, on a scale from 1 to 10.
c_class Binary color class for harvesting (Sufficient, Insufficient). Computed from c_val with threshold $6$.
l_val, r_val Expert given scores for lignification and berry shape development, respectively, on a scale from 1 to 10.
Table 2: These are the parameters considered by the agronomist in the exploratory experiment to assess the correlation between °Brix and human color evaluation.
In order to have an initial human performance baseline and to assess the experts’ actual capacity to estimate ripeness from appearance, we performed a specific test. We gave to a number of agronomists with experience in table grape ripeness estimation, $100$ grape images, for which a °Brix ground truth was collected with the technique described in Fig. 2. We used the JPG smartphone images collected during the first data collection session (24 September 2021) for Pizzutello Nero. The images were given without any other hint to the experts, asking them to express a numeric evaluation, on a scale from 1 to 10, of the degree of color, peduncle lignification stage, and berry shape. Note that this color scale is not the same as the one defined in Section 2.1.2. The reason is that this experiment uses a human-friendly decimal scale, while the one defined previously is to be used by a classification algorithm. In addition to the aforementioned parameters, the agronomists were asked to guess if the °Brix for the sample was high enough for harvesting purposes (in this case, $\geq 17$). Table 2 summarizes the experiment variables name and meaning. First, we checked the correlation between human-given values of c_val, l_val and r_val and the measured mean °Brix value, m_brix to explore the hypothesis where color and other visual elements could be a good predictor for sugar content.
In Fig. 6 the correlation matrix is shown. The correlation score between the average °Brix (m_brix) and color evaluation (c_val) is $0.505$, with $p<0.001$ and $CI=[0.342,0.638]$, the correlation between m_brix and lignification (l_val) is $0.337$, $p<0.001$ and $CI=[0.151,0.501]$, while the correlation with berry shape (r_val) score is much lower, $-0.038$ way below any significance threshold.
The observed correlation between human expert visual assessment and measured °Brix values suggests that a data-driven algorithm could achieve human-level accuracy in estimating grape quality for harvesting decisions. Even if there is a significant correlation with the lignification of the peduncle, in this work we focused on the color, leaving the use of peduncle lignification as a predictor for future work.
The second outcome of this experiment was to define some human harvesting performance baselines. For this question, we consider the c_class, the color category given by the agronomists to each sample, as the harvesting prediction and the thresholded values b_class as the binary labels. The precision and recall values are respectively $0.43$ and $0.83$, which means a false discovery rate (type I error rate) of $0.56$ and false negative rate (type II error rate) of $0.17$.
It is important to state that in the case of harvesting, type I errors (false positives) have a higher cost than type II since all the ripe bunches left on the vine tree will be eventually harvested, while the harvested unripe bunches have to be disposed of. We see that by simply collecting the table grape bunches with a color class over the threshold of $6$ (the definition for the c_class attribute) a lot of unripe bunches are collected due to °Brix to color mismatch. One question that arises then is if by increasing the threshold for c_class, it would be possible to trade type I errors for type II errors. Computing the precision and recall scores for a c_class value obtained with a threshold of $7$ for example, we get respectively $0.35$ precision and $0.37$ recall, which is an increase in both error types. This test shows that, while °Brix and color are correlated on the broad range of the color scale of this experiment, at the same time, the relationship is not quite linear, i.e. there are both a considerable number of insufficiently colored bunches that have sufficient °Brix, and highly colored bunches with insufficient °Brix. The high number of bunches on a borderline situation that the experts miss-classify shows that the question of training an appearance-based estimator is a non-trivial one.
Figure 6: Correlation matrix of the parameters estimated by human experts looking to a sample of 50 images of table grape bunches, as described in Section 3.1. The key factor for our evaluation is the correlation between the SSC guess (m_brix) and the evaluation of bunch color (c_val).
3.1.2 Color Constancy effect on Human Evaluation of Color
One of the early questions that arose from data collection was to what extent the camera image processing pipeline could influence the possibility of estimating the correct color stage, and therefore influence the °Brix estimate. Smartphone cameras, like most camera devices, do not process the raw signal from their CCD/CMOS sensors for radiation accuracy, but more for aesthetic purposes. This means that color accuracy is often lost in the nonlinear pipeline that transforms raw images into JPGs. A second consideration was related to the difference between the human vision system and a camera device. We know that human experts and workers alike are able to guess ripeness from appearance, but to what extent some of the peculiarities of the human vision system influence this guess (e.g. consider the brightness constancy effect) is unknown. While color perception is a complex topic, we focused our analysis on a specific question: whether human experts assign significantly different color labels when viewing white-balanced versus standard images. We selected 50 images from a single field trip that had paired raw and JPG image formats. The raw image was processed only with debayering and white balancing using the grey card. (Figure 7 shows some examples).
Figure 7: Some examples of pictures of the same table grape bunches were presented to agronomists on different occasions to test whether human experts consistently evaluated bunch ripeness, regardless of color correction, as described in Section 3.1.2
We presented all 100 images in random sequence to a group of agronomist experts, without revealing that each bunch appeared twice. The experts labeled the color quality of each image using the color scale described in Section 3, without performing any other manipulation or consideration. The null hypothesis is that there is no difference in labeling standard and white-balanced images apart from some generic labeling noise. If we define the mean group labeling error as:
$\mu_{i}=\mathbf{E}\left[L_{GT_{i}}-L_{exp_{i}}\right]$ (7)
where $i\in[0,1]$ is the sample group, $L_{GT_{i}}$ is the hypothetical ground truth color label for a random sample from group $i$, and $L_{exp_{i}}$ is the label that an expert gives for the same random sample. Furthermore, we denote with $Y_{i}$ the sample mean of the difference between the true label and the estimated one. Then, we can state that $Y_{1}-Y_{0}$ is an unbiased normally distributed estimator of $\mu_{1}-\mu_{0}$, and thus we can formulate the following null hypothesis:
$\mathcal{H}_{0}:\mu_{1}-\mu_{0}=0$ (8)
which we can easily test for significance with a z-test.
Analysis of the results revealed only a small difference between expert labels for the two groups, with a mean difference of $0.1$ and a standard deviation of $0.50$. The paired test for the sample batch gives a z-statistic of 1.4 (p=0.17), which does not provide sufficient evidence to reject $\mathcal{H}_{0}$. While a larger sample size might yield more definitive results, we chose instead to focus on the estimation algorithms, which are the main subject of our study. In particular, we included the comparison between white-balanced images and standard images into the cross-validation of the histogram-based features to gain deeper insights.
3.2 Histogram based color and °Brix estimation
This Section first outlines the experimental setup used to evaluate histogram features and subsequently presents the results.
3.2.1 Description
Given the exploratory experiments described in the previous Sections 3.1.1 and 3.1.2, we designed the histogram features as described in Section 2.2 and use them to train regression and classification models to predict SSC and color level, respectively.
The following experiments were conducted for both color estimators (multi-class classification) and SSC estimators (°Brix regression). Given the higher difficulty of SSC estimation compared to exterior color evaluation, the following description focuses on SSC estimation, while the tables and figures summarize the results for both SSC and color estimation.
The main question addressed in these experiments is the overall performance of the histogram-based algorithms and the influence of hyper-parameters. A number of grid hyper-parameter searches using Ridge Regression were conducted on a limited dataset, specifically the $T1_{STD}$ and $T1_{WB}$ splits, to mitigate distribution shifts due to different camera sensors and seasons during the initial analysis. These splits contain 112 samples each, divided into 89 train images and 23 validation images. The training images were augmented with horizontal flips, resulting in a total of 178 training images. As previously mentioned, both standard JPG and white-balanced images were evaluated to assess the influence of post-processing on the feature-based algorithms.
To set a baseline for this approach, we considered the intrinsic variability training and test data. The $T1$ datasets have a mean of 17.19 °Brix, a standard deviation of 2.43 °Brix, and a mean absolute deviation (MAD) of 1.99 °Brix. Appendix A provides more details on the statistical properties of the datasets. The naive predictor, which always predicts the mean °Brix values of the training set, resulting in a mean absolute error (MAE) equal to the MAD, serves as the lower bound for estimator performance.
Figure 8: a) Correlation between validation error and hyper-parameters for the top 10% performing runs. This plot highlights the direction (positive or negative) of the correlation. b) Test error for models with hyper-parameters identical to the best model shown in Table 3, but with varying $alpha$ values and color correction. The consistent lower error for models trained on standard images compared to color-corrected images gives a clear answer to the question raised in Section 3.1.2.
Test MAE Val MAE Train MAE $alpha$ $n\_bin\_x$ $n\_bin\_y$ $cross$ $color\_space$ $feature\_len$ $log(alpha)\times feature\_length$
1.459898 2.098344 1.583124 9.000000 16 8 none hsv 128 122.143041
1.509351 2.129067 1.548176 9.000000 8 16 none hsv 128 122.143041
1.522208 2.141015 1.614265 3.000000 8 8 none hsv 64 30.535760
1.509816 2.144949 1.178387 9.000000 8 32 none hsv 256 244.286082
1.507421 2.164503 1.619756 1.000000 16 8 thin lab 80 0.000000
1.495739 2.164640 1.254232 3.000000 16 8 none hsv 128 61.071521
1.517718 2.181812 1.431866 3.000000 16 8 fat hsv 96 45.803640
1.523156 2.193494 1.719893 0.330000 8 8 thin lab 40 -19.259442
1.431197 2.200048 1.189874 3.000000 8 16 none hsv 128 61.071521
1.491955 2.327332 1.225366 1.000000 16 8 thin hsv 80 0.000000
Table 3: The top 10 histogram-based models, ranked by validation mean absolute error (Val MAE), along with their corresponding hyper-parameters values. Some trends discussed in Section 3.2, such as the importance of alternative color spaces like HSV, are clearly evident.
3.2.2 Discussion
A grid search on the aforementioned hyper-parameter ranges, requiring training 3888 different models, was performed using the Weights & Biases experiment tracking framework [2] to compare and analyze the results. Further details of the hyper-parameter search are provided in Appendix B, while this section focuses on the key findings. The correlation coefficient of each feature with the validation error has been computed and is shown in Fig. 8(a). In addition, Table 3 shows the 10 best runs for the $low\_res$ images.
As expected, given the limited number of samples in the $T1$ datasets, the regularization parameters and those influencing feature vector length had the greatest impact on the final results. Furthermore, image resolution proved irrelevant, suggesting the safe use of lower-resolution images for computational efficiency in embedded applications. This is consistent with the feature construction process, where binning and averaging naturally remove the complexities of high-level details. Finally, and most notably, both Fig.8(a) and Table 3, indicate that using an alternative color space, such as HSV, helps in finding better coefficients for the Ridge Regression.
Focusing solely on the 1296 lower-resolution runs, the best-performing models have validation MAE of $1.45$, with different combinations of hyper-parameters that have in common a low feature count relative to the regularization value. Table 3 summarizes some of these top-performing runs, all exhibiting MAE lower than the baseline brixmad of 1.99. The best run reaches a $\mathbf{27\%}$ improvement over the baseline, a notable result considering the limited training data. Figure 8(b) displays the validation errors for different $alpha$ values (on a logarithmic scale), for both color-corrected and standard images. The most interesting feature of the plot, which is a common trend in all similarly produced plots, is the consistently inferior performance of the models trained on color-corrected images compared to their standard counterparts with identical hyper-parameters.
3.3 Multi-task Learning DNN
The SSC and color distribution can be viewed as conditioned on the environmental and agronomic parameters of the field we are considering. While both the color and the SSC depend on the internal chemical state of the berries, this state itself is influenced by environmental factors such as growing conditions, training practices, and nutrient availability, The overall appearance of the grape bunches is also affected by these factors, which we assume are homogeneous across the vineyard.
The architecture of the multi-task network, illustrated in Fig. 5, reflects this reasoning. The encoder component captures common features in appearance variations that may correlate with the internal chemical state of the grapes. These shared features are then leveraged by two separate prediction heads to estimate the color class and SSC in parallel. As before, in the following we first give a description of the experiment and then discuss the results.
3.3.1 Description
The $T2$ dataset (Table 1) is used to evaluate the covariate shift due to the variability in treatments between consecutive seasons while keeping the sensor type fixed. Since phone images were captured in both raw and JPG formats, the $T2$ splits include standard ($T2_{STD}$) and white-balanced ($T2_{WB}$) versions. This allows for a direct comparison of performances between standard JPG and color-corrected images, under identical hyper-parameters and for the same grape bunch instances. In both cases, data from the 2021 season serves as the training and validation sets, while the data from the 2022 season are used as the test set. The $T2_{ALL}$ split consists only of color-corrected images since it includes reflex camera images that have been captured in raw format and subsequently color-corrected using the gray card. This split helps assess to some degree the effect of using two different camera devices to capture the same instances with the same perspective. The $T3$ splits combine images from all devices across both seasons, including both color-corrected and standard versions.
Appendix C and Table 10 summarize the hyper-parameter ranges considered for the experimental evaluation. We performed $4912$ runs on a DGX server with four $A100$ GPUs. Each run was trained for $200$ epochs with hyper-band early stopping. For all the experiments we used a ResNet18 backbone [7], exploring both the effect of pretraining on ImageNet [4] and training from scratch. Hyper-parameter relevance was assessed using the BCH metric described in Section 2.3.1. While the complete details on the analysis of the importance of each hyper-parameter are described in Appendix C, in the following Section we discuss the most relevant findings.
Figure 9: Multi-task network hyper-parameter correlation - These histograms shows the correlation between the validation error and the hyper-parameters computed on the 25% best performing runs. a) $T2$ dataset: it can be noted that the main influence comes from the classical parameters that impact any DNN training (learning rate, batch size, regularization parameters), though image size and color space remain relevant. More insights on this can be extracted from Table 4. b) $T3$ dataset: similar trends are observed as in (a), with classical DNN training parameters having the most significant impact. The increased importance of blur parameters suggests the potential benefits of data augmentation strategies in mitigating the effects of covariate shifts due to variations in camera devices.
3.3.2 Discussion
We begin our analysis with the $1002$ runs performed on the $T2$ splits. Since we sampled a wide range of models via random search, we concentrate on the best $25\%$ of the total runs to discuss the influence of hyper-parameters. Figure 9(a) presents the correlation between the hyper-parameters and the BCH metric, while Table4 lists the top $10$ performing runs. The multi-task network effectively learns good models for both SSC and color estimation. Apart from the typical learning parameters, the ones that show more influence on the performance results are image size (particularly width) and color space (rarely RGB among top performers). However, no single hyper-parameter shows high correlation values, indicating a complex relationship linking the parameters to the error. Another consideration is that $\alpha$ is typically ten times larger than $\beta$, giving a rough estimate of the relative weight of the corresponding losses for the final results. Moreover, the benefit of the additional augmentations appears less substantial than initially anticipated. It is important to note that the lowest MAE for SSC and the highest accuracy for color estimation are not achieved at the same epoch. This is unsurprising, as perfectly correlated feature learning for distinct tasks is unlikely, even with strong underlying correlations. Therefore, when using a single feature extractor, a trade-off must be made between optimizing for SSC or color estimation. If computational resources permit, employing two different feature extractors (i.e., different checkpoints) allows for independent optimization and potentially better performance on both tasks.
Val BCH Val SSC MAE Val Acc $alpha$ $beta$ $im\_width$ $im\_height$ $momentum$ $learning\_rate$ $batch\_size$ $blur\_\sigma\_min$ $blur\_\sigma\_max$ $pretrained$ $one\_cycle$ $HSV$ $Lab$ $augment\_extra$
4.076824 1.555511 0.673913 500 50 256 256 0.990000 0.000001 64 0.100000 2 False False False True True
4.188802 1.832541 0.663043 1000 1000 256 64 0.950000 0.000001 32 0.500000 4 False True False True False
4.265176 1.653810 0.554348 100 10 64 256 0.950000 0.000050 64 0.500000 2 True False True False False
4.312880 1.732056 0.619565 500 100 256 64 0.800000 0.000050 128 0.100000 2 True True True False True
4.346299 1.579317 0.630435 1000 100 256 128 0.800000 0.000050 128 0.010000 4 False True True False False
4.412035 2.160111 0.532609 500 100 64 128 0.990000 0.000050 128 0.010000 2 False True True False False
4.453714 1.795929 0.684783 1000 500 256 256 0.950000 0.000010 128 0.010000 2 False False False True True
4.496901 1.891276 0.673913 100 50 256 128 0.900000 0.000050 128 0.500000 2 True True True False True
4.536293 1.629490 0.521739 100 100 256 128 0.900000 0.000050 64 0.500000 2 False False False False False
4.598450 1.620318 0.695652 10 1 256 256 0.990000 0.000050 128 0.100000 4 False True False True True
Table 4: The best ten runs on $T2$ split.
Similar trends emerge from the experiments on the $T3$ splits. Figure 9(b) shows the correlations of the hyper-parameters with the error, and Table 13 displays the best-performing runs. Again, the $\alpha$ to $\beta$ ratio is confirmed, along with the general preference for having higher image width values. The $T3$ test set, captured with a Realsense d435i, introduces a covariate shift due to its lower resolution and different white balance settings compared to the phone and reflex cameras used for training. This difference might explain the increased relevance of blur-related parameters in the augmentation.
Following the analysis conducted for histogram features in Section 3.2, we compare the best $T3$ run with runs using the same hyper-parameters but trained on white-balanced images. Unlike the histogram feature case with a single regularization parameter, the multi-task model utilizes two ($\alpha$ and $\beta$). Given what was observed from the experiments, that the ratio of the $\alpha$ and $\beta$ values is more important than their absolute value, we use this as a parameter to compare the impact of using color-corrected images versus using the standard JPG ones. Similarly to the analysis presented in Section 3.2, we select the 150 best-performing runs and the ones with matching hyper-parameters (except for $\log(\alpha/\beta)$) and use them to extract a general trend shown in the scatter plot in Figure 10. The regression lines for both standard and white-balanced image sets exhibit nearly identical slopes ($-0.212$ and $-0.216$, respectively), and the difference between the lines is small, with largely overlapping confidence intervals. This suggests that the multi-task model effectively extracts relevant features regardless of color correction, with performance oscillations likely stemming from training data variability and other hyper-parameters.
To conclude, the best-performing runs for SSC estimation achieve MAE values as low as 1.55 (1.57) for the model trained on the $T2_{STD}$ ($T2_{WB})$ split, and 1.11 (1.10) for the one trained on the $T3_{STD}$ ($T3_{WB})$ split. Considering that the MAD values for the $T2$ and $T3$ test splits are 1.71 and 1.36, respectively, the performance gain of the multi-task models is $9.3\%$ ($8.1\%$) for the $T2$ case and $18.4\%$ ($19.1\%$) for $T3$ case. These results confirm the negligible impact of color correction on the multi-task model performance, as observed in the regression analysis. The superior performance of the white-balanced model on the $T3$ dataset can be attributed to the automatic white balancing performed by the d435i camera used for collecting the $T3$ test set.
Figure 10: Linear regression analysis of runs with hyper-parameters similar to the top 10 runs on the $T3_{STD}$ and $T3_{WB}$ datasets. The shaded area represents the $95\%$ confidence interval. While the models trained on color-corrected samples present slightly higher values on average, the smaller difference compared to Fig. 8(b) suggests no practical advantage in using standard images over the white-balanced ones in this context.
4 Conclusions
This work addressed the problem of estimating soluble solid content (SSC) in grapes using RGB data acquired in the field. While numerous studies explore laboratory equipment and hyperspectral data for SSC estimation through visual sensors, this study is among the few to focus exclusively on the practical scenario of RGB data collected directly in field conditions. The complexity of this task arises from covariate shifts inherent to the problem, including seasonal variations and differences in imaging devices.
To tackle this challenge, a dataset spanning multiple years was collected, incorporating data from various camera devices and annotated for SSC and color quality based on agronomic practices. Human performance in assessing the correlation between color and SSC was evaluated through two user studies, which informed the design of data-driven algorithms for estimating both SSC and color.
Two algorithmic approaches were explored: a light-weight method using classical handcrafted features, suitable for devices with limited computational power, and a multi-task neural network architecture that, while more computationally demanding, effectively leverages correlations between visual appearance (color and shades) and SSC. Both methods were experimentally validated using field-collected data from the Pizzutello Nero grape variety. Results indicate that the proposed techniques achieve reliable SSC estimation with a reasonable margin compared to the average human estimator. Furthermore, experiments demonstrated some robustness to covariate shifts, including seasonal variations and device differences, positioning the proposed solutions as viable prototypes for low-cost, on-field applications where frequent data collection and labeling are challenging.
We release all the data, code, trained models, and training scripts used in this study. Future work will focus on refining and aligning these low-cost estimators and benchmarking them against hyperspectral approaches.
Acknowledgments
[…]
This work is part of a project that has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No.
101016906 – Project CANOPIES
[…]
This work has been partially supported by project AGRITECH Spoke 9 - Codice progetto MUR: AGRITECH ”National Research Centre for Agricultural Technologies” - CUP CN00000022, of the National Recovery and Resilience Plan (PNRR) financed by the European Union ”Next Generation EU”.
This work has been partially supported by Sapienza University of Rome as part of the work for project H&M: Hyperspectral and Multispectral Fruit Sugar Content Estimation for Robot Harvesting Operations in Difficult Environments, Del. SA n.36/2022.
Appendix A Detailed dataset description
Data collection was part of a larger effort within the Canopies project, exploring different sensors across seasons. The first data collection was done in September-October 2021, using a smartphone camera and a Reflex camera. The second data collection took place in September 2022, utilizing a smartphone camera and a Realsense d435i. Although the d435i is an RGB-D device, we used it as a standard camera for °Brix and color estimation.
Table 5 summarizes the different types of data collected during the two seasons. The naming of the splits in the following tables follows a descriptive naming used during the development of the dataset and the code. For ease of reference, we renamed the splits with simpler tags ($T1$, $T2$, $T3$) to avoid bloated experiment descriptions. The match between the splits names given in Table 1 and 7 are the following:
[…]
The data organization allows for testing covariate shifts due to both sensor type and season. These atomic splits represent minimal groups sharing the same sensor, season, image type, and label characteristics. Using these atomic splits, we designed the experimental splits, summarized in Table 6 and Table 7.
Images: Smartphone images were acquired at a resolution of 3000x4000 pixels, in both JPG and raw formats, with a few exceptions. Reflex images, captured in raw format, have a resolution of 3900x2613 pixels. The d435i RGB images have a lower resolution of 1280x720 pixels. Raw images were corrected for white balance using a gray card (Fig.2). Not all images included a gray card due to collection errors, but these images were still used for comparison purposes. The full images were labeled with bounding boxes around the grape bunches and then cropped. The resulting cropped images have very different sizes since some images were close-ups while others contained multiple targets. The combination of season, sensor used, gray card presence, and other factors creates differences summarized in Table 5. The rows in the table represent atomic splits, i.e., splits with the same distribution. In contrast, between the splits, there is some form of covariate shift that we aim to study.
Datset Id Year Smartphone Reflex d435i Num images Num samples Expert GT Color GT
Type Jpg Raw Card Raw Card RGB IR Card
A 2021 motog8 yes no no no - no no - 100 100 no no
B 2021 motog8 no yes yes yes yes no no - 38 38 possible yes
C 2021 motog8 yes yes yes yes yes no no - 62 62 possible yes
D 2021 motog8 yes yes yes yes yes no no - 50 50 possible yes
E 2022 realmi yes yes no no - yes no no 1 2 no no
F 2022 realmi yes yes yes no - yes no yes 6 22 possible yes
G 2022 realmi yes yes no no - yes no yes 1 2 no no
H 2022 realmi yes yes yes no - yes no yes 4 19 possible yes
I 2022 motog8 yes yes yes no - yes yes yes 8 51 possible yes
Table 5: Description of the atomic splits according to the device, image format, gray card presence, and the possibility of having expert ground truth from white-balanced images.
Name Components Description
2021PhoneJPG A+C+D Contains all the jpg images collected with a smartphone in 2021, without white balance from the spits A, C, and D.
2021PhoneJPG_reduced C+D Contains the jpg images collected with a smartphone in 2021 that have a corresponding white-balanced image. Corresponds to splits C and D.
2021PhoneJPG_WB_reduced C+D Contain all the white-balanced images collected with a smartphone in 2021. Corresponds to splits C and D.
2021PhoneJPG_WB B+C+D Contains all the jpg images collected with a smartphone in 2021, from splits B, C, and D
2021ReflexJPG_WB B+C+D Contains all the jpg images collected with a reflex camera in 2021 from splits B, C, and D
2022PhoneJPG_WB F+H+I Contains all the jpg images collected with a smartphone in 2022 from splits F, H, and I
2022d435iAutoWB F+H+I Contains all the auto white balanced jpg images collected with d435i in 2022, from splits F, H, and I
Table 6: Description of the basic splits used to train and test the proposed algorithms.
Name Train set components Test set components Description
Season21ToSeason22Phone A+C+D F+H+I All jpg images taken with a phone.
Season21ToSeason22Phone_WB B+C+D F+H+I All raw images taken with a phone that have been white-balanced.
Season21ToSeason22All_WB B+C+D F+H+I All phone and reflex images white-balanced.
Phone_d435i_Std A+C+D+F+H+I All phone images not white-balanced, plus all d435i images
AllWB_d435i_Std B+C+D+F+H+I All phone and reflex images white-balanced, plus all d435i images
MixedSeason_JPG A+C+D+H+I F All jpg images, keeping for testing only the ones with the most uncertain brix value (around 16).
MixedSeason_WB_WB B+C+D+H+I F All white balanced images, keeping for test only the most uncertain brix value (around 16).
Table 7: This table describes the most complex data splits, which we called mixed splits. These splits span across different sensors, while the train and test separation is based on the season.
Split name #$\#samples$ $Brix_{mean}$ $Brix_{std}$ $Brix_{mae}$ $CC_{1}$ $CC_{2}$ $CC_{3}$ $CC_{4}$ $CC_{5}$ $CC_{6}$
2021PhoneJPG 212 16.23 2.52 2.02 0.01 0.04 0.09 0.14 0.27 0.45
2021PhoneJPG_WB 150 17.54 2.46 2.04 0.01 0.03 0.06 0.10 0.21 0.59
2021PhoneJPG_reduced 112 17.19 2.43 1.99 0.01 0.04 0.07 0.12 0.22 0.54
2021PhoneJPG_WB_reduced 112 17.19 2.43 1.99 0.01 0.04 0.07 0.12 0.22 0.54
2022PhoneJPG 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
2022PhoneJPG_WB 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
2021ReflexJPG_WB 150 17.54 2.46 2.04 0.01 0.03 0.06 0.10 0.21 0.59
2022d435iAutoWB 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
MixedSeason_train 352 17.67 2.75 2.33 0.01 0.04 0.12 0.16 0.32 0.34
MixedSeason_test 44 15.96 1.68 1.36 0.00 0.00 0.09 0.09 0.27 0.55
MixedSeason_WB_train 440 18.27 2.40 1.98 0.01 0.04 0.10 0.13 0.27 0.46
MixedSeason_WB_test 44 15.96 1.68 1.36 0.00 0.00 0.09 0.09 0.27 0.55
Season21ToSeason22Phone_train 212 16.23 2.52 2.02 0.01 0.04 0.09 0.14 0.27 0.45
Season21ToSeason22Phone_test 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
Season21ToSeason22Phone_WB_train 150 17.54 2.46 2.04 0.01 0.03 0.06 0.10 0.21 0.59
Season21ToSeason22Phone_WB_test 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
Season21ToSeason22All_WB_train 300 17.54 2.46 2.04 0.01 0.03 0.06 0.10 0.21 0.59
Season21ToSeason22All_WB_test 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
All_Std 396 17.48 2.71 2.30 0.01 0.04 0.12 0.15 0.32 0.37
All_Std_WB 484 18.06 2.44 2.03 0.01 0.03 0.10 0.12 0.27 0.47
Table 8: Basic statistics for all the atomic and mixed splits described in Table 7. For each split the following statistics are listed: number of samples, average °Brix, standard deviation of the °Brix, °Brix mean absolute error, and the percentage of each color class in the dataset. The color classes are unbalanced towards the unripe color classes, since the ripe bunches were harvested every few days.
Appendix B Hyper-parameters Exploration for the Histogram based Estimators
Following from Section 3.2, here we give the details of the hyper-parameter search for the histogram features.
Table 9 summarizes the hyper-parameter ranges investigated by this experiment. Most of them have been introduced in Section 2.2. In addition to those, here we add the investigation of the effect of the input image resolution (low_res: $320\times 240$, med_res: $640\times 480$, high_res: $1280\times 720$). For the $cross$ parameter, to avoid hyper-parameter explosion, we designed three possible configurations (none, thin, fat), as described in Section 2.2.
Hyper-parameter Range description
$n_{bin_{x}}$ 4, 8, 16 number of image columns bins.
$n_{bin_{y}}$ 8, 16, 32 number of image rows bins.
$cross$ ”none”, ”thin” or ”fat” How many bins to drop for background elimination. See Eq.1 for details.
$color\_space$ ”none”, ”hsv” or ”Lab” Use of CIE-La*b*, HSV or RGB color coordinates.
$dataset$ low_res, med_res, high_res, Image resolution, corresponding to $320\times 240$, $640\times 480$ and $1280\times 720$ respectively, in color corrected and standard form.
wb_low_res, wb_med_res, wb_high_res
$\alpha$ 0.33, 1, 3, 9, 27, 81, 243, 729 Ridge Regression regularization parameter.
Table 9: Hyper-parameters ranges for the histogram feature based models.
To identify the most relevant hyper-parameters, taking into account the non-linear interactions between the different features, we opted to train a random forest (RF) algorithm on the hyper-parameter values themselves, to predict the validation error. Using this technique allows for reading the importance score of the RF as an indicator of the relevance of each hyper-parameter [5]. Figure 11(a) shows the importance factors for the RF regressor trained to predict the validation error from the hyper-parameter values. The importance factor show which features contributes the most to the RF regression of the validation error, allowing for concluding that the different starting image resolutions have little influence on the SSC estimation. This is a relevant finding, since it allow us to safely work with the lowest resolution images without trading for estimation accuracy. Among the most important variables we find the ones that control the model flexibility, i.e. the regularization parameter, the $x$ and $y$ bin number, and the cross parameters, that impact the feature vector length. Together these parameters have to balance prediction power with the risk of over-fitting.
Finally, the color space show some relevance too. The correlation coefficient has already been shown in Fig. 8(a), and its replicated in Fig. 11(b) for convenience. Image resolution was omitted and the total histogram length ($feature\_len$) was added. That histogram helps understanding which variable linearly correlates with the validation error, in positive or negative way, but alone it’s not enough to show feature importance, since there could be some interactions effects between them.
Figure 11: a) Training a Random Forest using the hyper-parameters values to predict the validation error and then extracting the importance scores is a common way to extract the relative impact of each feature on the final estimator performance. Here we can see that the choice of initial image resolution is not relevant, while the regularization, the number of bins (that influence feature length and flexibility), and the color space have to carefully selected. b) shows the correlation computed on the 10% best performing runs. This plot complements the other, showing where the correlation with the error is positive or negative.
Figure 12: To better extract the impact of feature length on validation error, we add three parameters to the analysis, $feature_{l}en$ and its interactions with $alpha$ and $\log{alpha}$. It is easy to see that the length of the histogram is the most important hyper-parameter to be tuned for final model performance.
To further explore this influence of feature flexibility due to histogram length, we show in Figure 12 the correlation between the validation error and the most important hyper-parameters extracted from Figure 11(a), explicitly adding $feature\_lenght$ and its interactions with $alpha$ and $\log(alpha)$. It is evident that the feature length is the most important factor among the others, and that the interaction between feature length and alpha is more than relevant. The general behavior that we can extract from this analysis is that the more are the bins of the histogram, the higher should be the alpha value, and together these two parameters are sufficient to explain a great deal of the validation error variability.
Appendix C Hyper-parameters Exploration for the Multi-task DNN
This appendix follow the same principles described in the previous one. Table 10 gives a summary of the hyper-parameters ranges explored in the sweep. Given the huge parameter space, this time the sweep was performed with random sampling, with a total run count of 4912 (see Section 3.3). Figures 13(a) and 14(a) show importance values for the RF trained on the hyper-parameter values to predict validation BCH respectively for $T2_{STD}$ (Season21ToSeason22_Phone) and $T3_{STD}$ (MixedSeason). Finally, Tables 4, 11, 13, 14, and 12 summarize the best 10 runs respectively on $T2_{STD}$, $T2_{WB}$, $T3_{STD}$, $T3_{WB}$ and $T2_{ALL}$.
Multi-task learning hyper-parameter Range description
$\alpha$ and $\beta$ 1, 10, 50, 100, 500, 1000 The loss balancing factors (Eq. 2)
$batch\_size$ 32, 64, 128 Training batch size
$learning\_rete$ 0.00001, 0.00005, 0.000001 Initial learning rate, to be modified by the scheduling strategy
$learning\_rete$ 0.00001, 0.00005, 0.000001 Initial learning rate, to be modified by the scheduling strategy
$momentum$ 0.8, 0.9, 0.95, 0.99 Momentum for the SGD optimizer
$im\_height$ and $im\_width$ 64, 128, 256 The image starting sizes before feature extraction by the auto-encoder.
$blur\_\sigma_{min}$ and $blur\_\sigma_{max}$ $\left[0.01,0.1,0.5\right]$ and $\left[2.0,4.0\right]$ Together they define the range for the Gaussian blur augmentation that is always applied.
$color\_space$ ”none”, ”hsv” or ”Lab” Use of CIE-La*b*, HSV or RGB color coordinates.
$one\_cycle$ $True$ or $False$ Whether to use of not the one cycle scheduling strategy
$pretrained$ $True$ or $False$ Whether the resnet18 backbone is pretrained on ImageNet or not.
$dataset\_split$ $Season21ToSeason22Phone$, $Season21ToSeason22Phone\_WB$, $Season21ToSeason22All\_WB$, $MixedSeason$, $MixedSeason\_WB$ Image resolution, corresponding to $320\times 240$, $640\times 480$ and $1280\times 720$ respectively, in color corrected and standard form.
$\alpha$ 0.33, 1, 3, 9, 27, 81, 243, 729 Ridge Regression regularization parameter.
Table 10: This Table summarizes the hyper-parameters descriptions for the Multi-task experiments. A wide range of hyper-parameters was explored using random search for a total of more than $5000$ experimental runs. In this case, the dataset splits used are not to be considered a proper hyper-parameter, but as a defining subset of the parameter sweep, since the distribution shifts between them are considerable. They are included in this table for ease of reference.
Figure 13: a) These plots show the importance and correlation factor of the best $25\%$ of the runs trained on the $Season2021ToSeason2022$ dataset split. It can be noted that the main influence comes from the classical parameters that impact any DNN training (learning rate, batch size, regularization parameters), while the choice of image size and color space comes seems to be still relevant for the final performances. More insights on this can be extracted from Table 4.
Figure 14: These plots show the importance and correlation factor of the best $25\%$ of the runs trained on the $MixedSeason$ dataset split.
Val BCH Val SSC MAE Val Acc $alpha$ $beta$ $im\_width$ $im\_height$ $momentum$ $learning\_rate$ $batch\_size$ $blur\_\sigma\_min$ $blur\_\sigma\_max$ $pretrained$ $one\_cycle$ $HSV$ $Lab$ $augment\_extra$
4.100699 1.835277 0.478261 500 50 128 256 0.800000 0.000050 64 0.500000 2 True True True False True
4.132786 1.913489 0.576087 100 1 256 64 0.990000 0.000050 64 0.010000 2 True True False False True
4.384027 1.570156 0.554348 1000 10 256 256 0.990000 0.000001 32 0.010000 2 True False False False True
4.397698 1.605315 0.434783 500 10 128 128 0.800000 0.000050 32 0.500000 2 False True False True True
4.499346 1.982949 0.576087 100 100 256 256 0.950000 0.000050 64 0.010000 4 True True True False False
4.522784 1.714363 0.500000 1000 100 256 128 0.990000 0.000010 64 0.100000 2 True False False True False
4.552584 1.768912 0.489130 500 10 128 128 0.900000 0.000050 64 0.500000 2 True True False True False
4.558001 1.644159 0.445652 1000 500 256 128 0.800000 0.000010 32 0.100000 4 False True False True False
4.583944 1.916520 0.521739 500 10 256 256 0.950000 0.000010 64 0.500000 4 False False False False True
4.596296 1.968404 0.597826 100 1 64 64 0.990000 0.000050 32 0.100000 4 True True True False True
Table 11: The best ten runs on $Season2021ToSeason2022\_WB$ split. Compared to the ones given in Table 4, both the SSC MAE estimation error and Color estimation accuracy show worse performance.
Val BCH Val SSC MAE Val Acc $alpha$ $beta$ $im\_width$ $im\_height$ $momentum$ $learning\_rate$ $batch\_size$ $blur\_\sigma\_min$ $blur\_\sigma\_max$ $pretrained$ $one\_cycle$ $HSV$ $Lab$ $augment\_extra$
3.977163 1.754222 0.576087 50 1 128 128 0.990000 0.000050 32 0.500000 2 True True False True True
4.123796 1.945782 0.532609 100 10 128 128 0.950000 0.000050 64 0.010000 2 False False True False False
4.269722 1.610763 0.673913 500 500 256 256 0.990000 0.000010 64 0.100000 4 True True True False True
4.307943 1.754948 0.608696 500 10 128 256 0.950000 0.000010 32 0.100000 2 True True False False False
4.327728 1.697267 0.467391 1000 50 128 128 0.950000 0.000010 32 0.500000 4 False True False True False
4.336076 1.930476 0.532609 1000 10 64 256 0.800000 0.000050 64 0.010000 2 True False False True False
4.347323 1.688515 0.500000 500 100 256 256 0.900000 0.000010 64 0.100000 4 False False False True True
4.411548 1.531855 0.597826 500 50 256 256 0.990000 0.000010 32 0.500000 2 False False False False False
4.440626 1.685545 0.630435 1000 500 256 256 0.990000 0.000001 128 0.100000 2 False False True False False
4.501309 1.803873 0.608696 1000 10 256 256 0.800000 0.000050 64 0.100000 4 True False False True True
Table 12: The best ten runs on $Season2021ToSeason2022All\_WB$ split.
Val BCH Val SSC MAE Val Acc $alpha$ $beta$ $im\_width$ $im\_height$ $momentum$ $learning\_rate$ $batch\_size$ $blur\_\sigma\_min$ $blur\_\sigma\_max$ $pretrained$ $one\_cycle$ $HSV$ $Lab$ $augment\_extra$
1.581554 1.200569 0.795455 500 10 128 64 0.900000 0.000050 32 0.010000 2 True False True False True
1.662561 1.116875 0.772727 1000 50 256 128 0.990000 0.000010 32 0.100000 4 True False False False True
1.686950 1.162343 0.772727 1000 50 256 256 0.950000 0.000010 32 0.010000 4 True True False True False
1.723328 1.231537 0.795455 50 100 256 256 0.950000 0.000010 32 0.010000 4 False False True False True
1.741897 1.102953 0.795455 1000 1000 64 64 0.990000 0.000001 32 0.500000 4 False True False True True
1.745264 1.301862 0.795455 1000 1000 128 128 0.950000 0.000001 32 0.100000 4 False True False False False
1.746910 1.269346 0.818182 1000 1000 256 256 0.900000 0.000010 128 0.100000 2 True True False True False
1.771968 1.174211 0.795455 100 500 64 128 0.950000 0.000010 64 0.100000 4 False False False False True
1.788137 1.190380 0.772727 100 50 128 256 0.950000 0.000050 128 0.500000 4 False False False False False
1.788171 1.242672 0.818182 10 10 256 256 0.950000 0.000050 64 0.100000 4 True True False False True
Table 13: The best ten runs on $MixedSeason$ ($T3_{STD}$) split. The validation set is the same as for the $MixedSeason\_WB$ ($T3_{WB}$), that is collected with a d435i camera. Since the images from this device are automatically white balanced, this imply that the model faces a strong covariate shift due to different camera device and image modality.
Val BCH Val SSC MAE Val Acc $alpha$ $beta$ $im\_width$ $im\_height$ $momentum$ $learning\_rate$ $batch\_size$ $blur\_\sigma\_min$ $blur\_\sigma\_max$ $pretrained$ $one\_cycle$ $HSV$ $Lab$ $augment\_extra$
1.545861 1.307749 0.818182 50 100 64 128 0.800000 0.000050 32 0.500000 4 True True False False False
1.596429 1.101236 0.795455 100 100 128 256 0.800000 0.000010 128 0.010000 2 False False True False False
1.611784 1.219757 0.772727 500 1000 128 128 0.800000 0.000010 32 0.010000 4 False False True False False
1.639119 1.065401 0.727273 50 10 128 256 0.950000 0.000050 64 0.500000 2 False True False False True
1.688067 1.220433 0.750000 500 500 128 128 0.950000 0.000010 32 0.500000 4 True False True False False
1.709684 1.295742 0.704545 10 10 256 256 0.990000 0.000010 32 0.010000 2 True False False True False
1.715017 1.159092 0.772727 100 100 64 128 0.800000 0.000050 32 0.100000 2 True False True False False
1.715871 1.158179 0.840909 10 10 256 256 0.900000 0.000050 64 0.100000 2 True False True False True
1.759702 1.200407 0.795455 50 50 128 128 0.900000 0.000050 128 0.100000 4 False False True False False
1.776299 1.200725 0.772727 100 50 256 256 0.900000 0.000050 32 0.010000 4 True True True False False
Table 14: The best ten runs on $MixedSeason\_WB$ ($T3_{WB}$)split. This time the results are comparable with the STD ones presented in Table 13. This can be explained considering the fact that the validation images of the d435i are white-balanced and therefore with a similar hue to the color-corrected training images.

### v2

@@
 This study investigates the potential of simple RGB sensors under uncontrolled lighting to estimate SSC and color, enabling cost-effective, robot-assisted harvesting.
-Over the 2021 and 2022 summer seasons, we collected grape images with corresponding SSC and color labels to evaluate algorithmic solutions for SSC estimation on embedded devices commonly used in robotics and smartphones.
-Our results demonstrate that SSC can be estimated from visual appearance with human-like performance.
-We propose computationally efficient histogram-based methods for resource-constrained robots and deep learning approaches for more complex applications.
+Over the 2021 and 2022 summer seasons, we collected grape images with corresponding SSC and color labels to evaluate algorithmic solutions for SSC estimation, specifically testing for cross-seasonal and cross-device robustness.
+We propose two approaches: a computationally efficient histogram-based method for resource-constrained robots and a Deep Neural Network (DNN) model for more complex applications.
+Our results demonstrate high performance, with the DNN model achieving a Mean Absolute Error (MAE) as low as $1.05$Brix on a challenging cross-device test set.
+The lightweight histogram-based method also proved effective, reaching an MAE of $1.46$Brix.
+These results are highly competitive with those from hyperspectral systems, which report errors in the 1.271.271.271.272.202.202.202.20 °Brix range in similar field applications.
 keywords:
@@
 […]
-[inst1]organization=Sapienza University of Rome Department of Computer, Control and Management Engineering ,city=Rome, postcode=00185, state=RM, country=Italy
+[…]
 1 Introduction
-The use of passive imaging sensors to measure agronomic properties of crops enables fast, non destructive monitoring that, when integrated with machinery or robots, can be efficiently scaled to cover entire fields.
-RGB cameras are commonly employed to monitor external conditions, such as detecting pests and diseases, while multi-spectral cameras are occasionally used to derive indices related to vegetation health.
+The use of passive imaging sensors to measure agronomic properties of crops enables fast, non-destructive monitoring that, when integrated with machinery or robots, can be efficiently scaled to cover entire fields (Fig.
+1).
+RGB (Red, Green, Blue) cameras are commonly employed to monitor external conditions, such as detecting pests and diseases, while multi-spectral cameras are occasionally used to derive indices related to vegetation health.
 However, assessing internal crop properties presents a more complex challenge.
-Hyperspectral imagery (HSI) show potential in this area, as their ability to capture specific wavelength responses can be correlated with the absorption patterns of certain chemical properties.
+Hyperspectral imagery (HSI) shows potential in this area, as its ability to capture specific wavelength responses can be correlated with the absorption patterns of certain chemical properties.
 Nevertheless, hyperspectral cameras are typically slow to acquire images, highly sensitive to illumination conditions, and usually unsuitable for use on operational agricultural machinery.
-Figure 1: Computer Vision plays a crucial role in enabling automatic harvesting operations; however, the requirements for high-quality performance often conflict with the economic and practical constraints of robot design.
-This work demonstrates the development and deployment of a simple RGB-based algorithm for SSC and color quality estimation, addressing the complexities of field operations while utilizing cost-effective sensors.
-These competing aspects collide with the desirable tension to bring more robots on the farming fields, in order to both face the shortage of workforce, but also to manage the crops in a more precise way.
+Figure 1: A robotic platform for grape harvesting developed within the CANOPIES project ([1]), which provides the context for this research.
+Such robots require robust perception systems to assess fruit quality in the field using on-board sensors The end-effector and the head of the robot are equipped with simple, low-cost RGB cameras suitable for this task.
+This work focuses on developing and validating the core data-driven algorithms necessary for SSC and color estimation in these challenging robotic setups.
+Table 1: Comparison of Hyperspectral (HSI) and RGB Imaging devices for in-field use.
+Feature Hyperspectral Imaging (HSI) Standard RGB Imaging
+Spectral Resolution High (hundreds of narrow bands, usually 5–10 nm) Low (3 broad bands: Red, Green, Blue)
+Spatial Dimension Low High
+Information Captured Detailed spectral signatures, correlated with chemical composition Surface color, texture, and shape
+Equipment Cost $10000$$200000+$ $10$$1000+$
+Operational Environment Laboratory or highly controlled settings In-field or indoors alike
+Acquisition Time High (seconds to minutes, depending on illumination) Low (few milliseconds)
+Computational Load High (due to the processing 3D hypercubes) Low (suitable for real-time processing on embedded systems)
+Portability & Size Often bulky and requires careful calibration Compact, and easily integrated into robotic platforms
+These competing aspects are at odds with the desirable tension to bring more robots to the farming fields, in order to both face the shortage of workforce, but also to manage the crops in a more precise way.
 In fact, many cultivations are labor-intensive, and for various economic and social reasons, it is increasingly difficult to find human workers to perform operations such as harvesting.
@@
 This shortage would be mitigated by the introduction of robot workers that could, to some extent, assist human workers in the field.
-To be able to do so, the robotic platform should be able to assess the quality and ripeness of the fruit using mainly passive sensors (active manipulation for perception would be orders of magnitude slower).
-While solutions to equip robots with hyperspectral cameras are becoming more prevalent, their high cost and operational complexity remain major barriers to widespread adoption.
-So, the question of designing a passive quality estimation system using low-cost, reliable, RGB cameras, or similar devices is still open.
+To be able to do so, the robotic platform should be able to assess the quality and ripeness of the fruit, such as color, Soluble Solid Content (SSC), size, presence of anomalies, using mainly passive sensors (active manipulation for perception would be orders of magnitude slower).
+The case of SSC estimation is of particular relevance as its estimation via passive sensors is usually done with Infrared (IR) refractometers (that require at least contact with the berries) or hyperspectral imaging (HSI) devices.
+Although solutions equipping robots with HSI cameras are becoming more prevalent, their high cost and operational complexity remain major barriers to widespread adoption compared to the use of other devices, such as RGB cameras (Table 1 offers a summary of the main differences between hyperspectral and RGB cameras).
+So, the question of designing a passive SSC estimation system using low-cost, reliable, RGB cameras, or similar devices is still open.
 This work addresses this question by validating its feasibility and proposing high-performance algorithmic solutions tailored for practical deployment in agricultural settings.
 In the following paragraphs, we provide a brief review of the relevant literature, highlighting key contributions to contextualize this study.
-For a broader overview of fruit ripeness estimation methods—particularly for grapes using hyperspectral and RGB imaging—we refer the reader to [11, 17].
-1.0.1 Hyperspectral applications to agricultural robotics
-Robots have been used for the monitoring and quality estimation of different crops and are now finding some industrial applications in high-value crops (wine grapes [20, 12], strawberries [23, 22], blueberries [19], among others).
+For a broader overview of fruit ripeness estimation methods—particularly for grapes using hyperspectral and RGB imaging—we refer the reader to Palumbo et al., [12] and Vrochidou et al., [20].
+Table 2: Comparison of different approaches for SSC estimation in Table Grapes using spectral equipment.
+Name Camera model Camera type Spectral Range Spectral Res. Spatial Res. Cultivar Illumination
+Gutiérrez et al., [6] Resonon Pika Push broom VIS-NIR (400-1000nm) 2.1nm (300 bands) 300 px Tempranillo Sunlight
+Kalopesa et al., [10] Contact probe spectrometer Spectrometer (contact probe) VNIR–SWIR (350–2500 nm) Not specified N/A (point measurement) Chardonnay, Malagouzia, Sauvignon-Blanc, Syrah Internal illumination
+Tsakiridis et al., [18] Cubert FireflEYE V185 Snapshot VIS-NIR (400–1000 nm) 4-8nm (128 bands) 50x50 Chardonnay, Malagouzia, Sauvignon-Blanc, Syrah Sunlight with learned calibration
+1.1 Hyperspectral applications to agricultural robotics
+Robots have been used for the monitoring and quality estimation of different crops and are now finding some industrial applications in high-value crops (e.g.
+wine grapes in Williams et al., [23] and Polvara et al., [13], strawberries in Xiong et al., [26, 25], blueberries in Williams et al., [22], among others).
 Some authors have explored specifically the quality estimation of fruits with hyperspectral sensors.
-In [6] an HSI device is mounted on an AGV to measure the TSS of Tempranillo wine grapes.
+In Table 2 we summarize the HSI approaches applied to grapes internal quality estimation.
+In Gutiérrez et al., [6] an HSI device is mounted on an Autonomous Ground Vehicle (AGV) to measure the Total Soluble Solids (TSS) — a synonym of SSC — of Tempranillo wine grapes.
 The AGV moves along the vineyard row acquiring spectral data in a push-broom fashion.
 To achieve color constancy the authors perform an indirect calibration for exposure, while illuminant calibration was performed every 5m with a white reference.
-The impact of these choices on the accuracy of the estimation is not clear, and the authors report an RMS on °Brix estimate as low as $1.274\textdegree$.
-In addition, the method requires considerable post-processing time (5 hours), making it not well suited for robotic harvesting.
-[18] propose a trailing cart mounting two HSI devices to monitor raspberries.
-However, the measurement is taken using an overarching arm from which a white reference is hung.
-[16] propose the Ladybug robot equipped with an HSI device built for low horticultural crops.
-The robot AGV is a straddle model, where the illumination constancy is granted through the use of a stroboscopic light.
-In preparation for robotic use, [9] perform a very thorough experimentation using VIS-SWIR images to estimate SSC in four wine grapes varieties.
+Although the authors claim that this setup shows on-the-go HSI estimation feasibility, they also note the difficulty of maintaining the steady motion required by push-broom cameras using an AGV.
+In addition, white balancing stop-and-go operations are required every 5 meters to calibrate the camera.
+The impact of these choices on the accuracy of the estimation is not clear, and the authors report a Root Mean Square Error (RMSE) on °Brix estimate as low as $1.274$°.
+Finally, the method requires considerable post-processing time (5 hours), making it not well suited for robotic harvesting.
+Being able to use a simple RGB camera for the estimation would remove both the vibration and calibration issues.
+A similar approach, but with a trailing cart instead of an AGV, was proposed by Williams et al., [21].
+The cart mounts two HSI devices (VNIR: 400-896nm, SWIR: 896-2506nm) to monitor raspberries.
+Again, the push broom image is obtained by moving a vehicle and the whole setup requires an overarching arm with a black background screen behind the plants and from which a white reference is hung.
+The aim of the study was to just demonstrate the setup and no physical characteristics have been estimated using the hyperspectral data.
+A very successful robotic platform, called Ladybug, is presented by Underwood et al., [19].
+The Ladybug robot, among other sensors and navigation equipment, is also equipped with an HSI device (Resonon Pika II, VNIR: 390-887nm) built for low horticultural crops.
+The robot AGV is a straddle model, where the illumination constancy is granted through the use of a stroboscopic light, in order to do not interfere with the rest of the optical sensors.
+In addition, calibration pads have been placed both on the field and on the robot, on the periphery of the HSI device’s field of view (FOV).
+The camera has been used to compute only the Normalized Difference Vegetation Index (NDVI), as the whole application was built to gather phenotyping information such as the plant height.
+The use of straddle systems is one of the most promising strategies for the application of HSI in the fields, however the calibration challenges remain, and the use for taller crops is unpractical.
+In preparation for robotic use, Kalopesa et al., [10] perform a very thorough experimentation using VIS-SWIR contact spectrometer (400-2500nm) to estimate SSC in four wine grapes varieties, namely Chardonnay, Malagouzia, Sauvignon-Blanc and Syrah.
+Contact probe spectrometers have a similar output to HSI imaging devices, but are meant for precise point sampling.
+These devices provide an internal illumination source and have to be carried around with a handle, or in a backpack.
+The cost is still higher than RGB cameras, and more complex to integrate into a robotic platform due to the delicate physical contact requirement.
 The authors analyze the contribution of different wavelengths to the estimation algorithms in depth and apply various algorithms for the prediction.
-Across the proposed models (Partial Last Square, Random Forest, Convolutional Neural Networks, and Support Vector Machines) and four grape varieties they obtain an RMS error ranging from 1.70 to 2.20 °Brix.
-In a follow-up study, to address the problem of illuminant estimation, [15] propose to learn the illuminant normalization trough the use of an auto-encoder.
-The solution is similar to ours, in the sense that it learns the typical illuminant distribution, but loses the absolute metric accuracy of having a white reference in the scene.
+Across the proposed models – Partial Least Square Regression (PLSR), Random Forest (RF), Convolutional Neural Networks (CNN), and Support Vector Machines (SVM) – and four grape varieties they obtain an RMSE ranging from 1.70 to 2.20 °Brix.
+In a follow-up study (Tsakiridis et al., [18]), in addition to the spectrometer, the authors introduce the in-field estimation of SSC with the use of a snapshot HSI camera (the Cubert FireflEYE V185, VNIR: 400-1000nm).
+In a different fashion from other works the main contribution of this work is a method to address the problem of illuminant estimation without the use of a reference tile before each acquisition.
+Instead they propose to learn the illuminant normalization through the use of an auto-encoder.
+The ground truth for the SSC estimation is collected with an analog refractometer, sampling the °Brix of a single berry for each bunch.
+The solution of using a deep neural network to estimate the illuminant distribution bears some similarities with our work as it simplifies the acquisition setup, but loses the absolute metric accuracy of having a white reference in the scene and still requires periodic re-calibration of the network using an more reliable spectrometer.
+In addition, the use of a snapshot HSI camera considerably reduces the spatial resolution, with respect to push broom setups.
 The authors explore four varieties of wine grapes, declaring RMSE values that range from 1.66 to 2.20 °Brix.
@@
 In contrast, the approach proposed in this work achieves comparable performance without relying on HSI, operates in real-time, and is thus better suited for integration into robotic systems.
-1.0.2 RGB applications to fruit quality assessment
-Given the limitations of HSI illuminant and color calibration and the cost of hyperspectral devices, many authors proposed approaches based on standard RGB sensors.
-[1] have explored the use of color imaging combined with advanced numerical techniques, such as interval numbers (INs), and neural network regressors to estimate grape maturity indices (e.g., TSS, TA, and pH), demonstrating the feasibility of non-destructive and cost-effective methods for integration into autonomous harvesting systems.
-In this work, the authors declare a considerable 6.6259 MSE Baumé, which approximately corresponds to an RMSE of 4.63 °Brix.
-[21] perform a very interesting analytical study on 180 samples of Kyoho grape berries using an RGB camera.
-The authors propose the use of some feature indexes, similar in principle to the well-known NDVI, but computed with the standard RGB channels, obtaining a very low cross-validation error of 0.78 °Brix (RMSE) on a dataset with a standard deviation of 1.37°.
-The study is performed in laboratory conditions, but it is nonetheless significant for our work, as it shows the effectiveness of data-driven estimations with simple RGB cameras.
-[3] proposed a non-destructive estimation technique based on an RGB camera using random forest predictors trained on color histograms.
+1.2 RGB applications to fruit quality assessment
+Given the limitations of HSI illuminant and color calibration, and the cost of hyperspectral devices, many authors proposed approaches based on standard RGB sensors.
+Bazinas et al., [2] have explored the use of color imaging combined with advanced numerical techniques, such as interval numbers (INs), and neural network regressors to estimate grape maturity indices (e.g., Total Soluble Solids (TSS), Titratable Acidity (TA), and pH), demonstrating the feasibility of non-destructive and cost-effective methods for integration into autonomous harvesting systems.
+The authors use a stereo camera (ZED Mini) to acquire RGB images converted to CIELab color space, under natural daylight.
+The grape variety is Tempranillo that, as in our case, is a black grape cultivar, that allows for an easier correlation of color and SSC.
+While similar to our approach, the declared performances are considerably higher, i.e.
+6.6259 Baumé Mean Squared Error (MSE), which approximately corresponds to an RMSE of 4.63 °Brix.
+Xia et al., [24] perform a very interesting analytical study on 180 samples of Kyoho grape berries using Samsung industrial RGB camera (SCC-B1011).
+The authors propose the use of some feature indices, similar in principle to the well-known NDVI, but computed with the standard RGB channels, as the inputs to two estimation methods, Partial Least Squares Regression (PLSR) and Multiple Linear Regression (MLR), obtaining a very low cross-validation error of 0.78 °Brix (RMSE) on a dataset with a standard deviation of 1.37°.
+The study is performed under laboratory conditions, but it is nonetheless significant for our work, as it shows the effectiveness of data-driven estimations with simple RGB cameras.
+In this work we show that also in uncontrolled environments it is possible to achieve estimation errors low enough to perform automatic harvesting reliably.
+Table 3: Summary of RGB-based approaches for fruit quality assessment (focus on SSC).
+Reference Camera Model Cultivar(s) Illumination Key Method / Features SSC Performance
+Bazinas et al., [2] ZED Mini 3D Tempranillo (100 samples of 10 vine trees in multiple dates) Daylight Color imaging, Interval Numbers (INs), Neural Network regressors RMSE 4.63 °Brix
+Xia et al., [24] Samsung SCC-B1011 camera Kyoho (180 berry samples) Laboratory conditions Custom RGB feature indices; PLSR and MLR RMSE 0.78 °Brix (Cross-Validation).
+Cavallo et al., [4] RGB camera Italia, Victoria (200 bunches) Artificial illumination CIELab color histograms (clustered, high variance colors), Random Forest, background subtraction.
+5-level quality scale.
+Cavallo et al., [4] proposed a non-destructive estimation technique based on an RGB camera (JAI CV-M9GE industrial CCD camera) using random forest predictors trained on color histograms.
 The histogram was computed in CIELab color space by clustering the colors and then choosing the ones with the highest intrinsic variance.
+While this is the only study that targets white grapes, the estimation problem was framed as a qualitative color scale measurement.
 The classification task was performed on a 5-level quality scale, and no explicit SSC measurement was included in the labeling process.
-The experimental setup involved 200 bunches, evenly divided between Italia and Victoria varieties, shot with artificial illumination, and image features were computed after background subtraction.
+However, the harvested bunches provided were nominally above 16° Brix level, that for the table grape food industry is a common quality requirement.
+The experimental setup involved 400 bunches, evenly divided between Italia and Victoria varieties, shot with artificial illumination, and image features were computed after background subtraction and careful white balancing using color calibration references.
+While this study cannot be quantitatively compared to ours, it is noteworthy as it shows that even with white grape varieties the correlation between color and SSC can be captured by careful modeling.
+The main limitation of this work that we are trying to overcome is the work in a controlled environment.
+For this reason, in this initial study we focus our attention to a black grape variety, as in Bazinas et al., [2] and Xia et al., [24].
+Table 3 summarizes these RGB-based approaches to SSC estimation in grapes.
 This growing body of research on RGB-based estimation of internal crop characteristics highlights both the scientific and practical interest in this approach.
 However, to the best of our knowledge, existing methods either fall short of achieving the performance levels of HSI imaging or fail to transition effectively from controlled laboratory conditions to field environments with uncontrolled lighting.
-Building on the challenges and opportunities highlighted in previous works, this study focuses on the problem of perception for quality estimation in table grape harvesting.
-Accurate quality estimation can enable robots to assist human workers or even surpass their capabilities during field operations.
-In high-quality table grape production, harvesting decisions typically follow a two-step process: an agronomist first samples grape bunches across the vineyard using a refractometer to measure SSC in °Brix.
-If the °Brix is sufficiently high, additional visual characteristics, such as color intensity and distribution, are evaluated.
-Workers are then instructed to harvest grape bunches based solely on appearance and location within the vineyard.
-While °Brix and color are not causally linked, they share latent ripening factors [8], allowing human experts to approximate °Brix from color to some extent.
-This work aims to replicate and enhance this human capability using data-driven methods and low-cost RGB sensors.
-The contributions of our study are as follows:
+This study addresses the challenge of perceptual quality estimation for robotic table grape harvesting, a task that could enable robots to assist or even surpass human workers.
+We model our approach on the standard harvesting procedure where an agronomist first confirms a target SSC in °Brix using a refractometer.
+Subsequently, workers harvest individual bunches based on visual cues like color.
+Although SSC and color are not causally linked, they share latent ripening factors – such as climate, weather, training practices and use of nitrogen, to name a few (Jackson, [9]) – which allows human experts to visually approximate ripeness.
+Our central goal is to replicate and enhance this expert capability using data-driven methods with low-cost RGB sensors.
+Our primary contributions are:
   • 1.
-    A dataset of table grape images collected across different seasons and with different sensors, labeled with both °Brix values and color categories following the aforementioned agronomic procedure, thus enabling studies on SSC estimation in realistic field conditions.
+    A multi-season, multi-sensor dataset of in-field grape images, labeled with both SSC (°Brix) and agronomic color categories.
+To the best of our knowledge, this is the first publicly available dataset that combines images with °Brix labeling, thus enabling studies on quality estimation under realistic field conditions.
   • 2.
-    Experiments and user studies to explore the extent to which °Brix can be estimated from color alone.
-These insights informed the design of algorithms and highlighted key factors for future research.
+    An analysis of the visual-SSC correlation, including a user study that establishes a baseline for human-level performance in estimating ripeness from images.
   • 3.
-    Two different estimation algorithms have been proposed and validated:
+    The design and validation of two distinct algorithms for real-time, calibration-free operation:
     1. a)
-      A lightweight method based on handcrafted features for constrained computational systems.
+      A lightweight, histogram-based method for resource-constrained robotic systems.
     2. b)
-      A deep, fully convolutional multi-task architecture that achieves high performance with limited training data.
-  • 4.
-    The methods do not require complex calibration and are designed for real-time use in robotic harvesting or mobile devices.
-The paper is structured as follows: Section 2 describes the dataset that was collected and the estimation methods that were designed; Section 3 defines the experimental setup, and describes and comments on the experimental validation of the problem and the algorithms.
-Finally, Section 4 draws conclusions and outlines a path for future work.
+      A deep, fully convolutional network that achieves high performance even with limited data.
+The remainder of this paper is structured as follows.
+Section 2 details our dataset and proposed estimation methods.
+Section 3 presents the experimental validation and and Section 4 discusses performance, limitations and computational constraints.
+Finally, Section 5 concludes the paper and outlines future research directions.
 2 Materials and Methods
+In this Section we detail the data acquisition and labeling processes and the two algorithmic approaches to SSC estimation. Fig, 2 shows the general overview of the methodology.
+Figure 2: An overview of the experimental workflow presented in this paper.
+The process moves from left to right: (1) In-field Data Collection using multiple RGB sensors and subsequent labeling for SSC and color quality.
+(2) Organization of the data into three distinct Datasets to evaluate performance under different conditions.
+(3) Development and validation of two estimation algorithms: a lightweight histogram-based model and a high-performance multi-task DNN.
+(4) The Test panel summarizes the key result, outlining our best model’s SSC estimation error (MAE) and harvesting decision performance (Precision/Recall) against the human expert baseline.
 2.1 Dataset
@@
 We collected static images during two summer seasons using different camera sensors.
-Three different devices, across multiple years, have been used to collect the images.
-We employed a smartphone, a Reflex, and a Realsense d435i to represent some of the possible covariate shifts that occur between the collection of the initial data on the field by a farmer or an agronomist (smartphone), a high-quality sensor (reflex) and a low-quality robotic camera (d435i).
-Table 1 depicts the working categories we used in this study.
-A more detailed description of all the data collected and the different characteristics is given in Appendix A for the interested reader.
+Three different devices, across multiple years, have been used to collect the images to represent some of the possible covariate shifts that occur between the collection of the initial data on the field by a farmer or an agronomist and the deployment on a robotic hardware.
+Images capture: We acquired images using two different smartphones, a Reflex camera, and a Realsense d435i.
+The smartphone images were captured with a Moto G8 Plus and a Realme GT Master Edition, respectively at $3000\times 4000$ and $3448\times 4624$ pixel resolution in JPG and raw formats.
+The Reflex images were captured in raw format at a $3900\times 2613$ pixel resolution, while the d435i provided RGB images at a lower $1280\times 720$ pixel resolution.
+When available, raw images were corrected for white balance using a gray card (Fig.
+3), though images without a gray card were still used for comparison.
+Fig.
+4 shows some examples of images collected with the same smartphone both in JPG and white-balanced RAW format.
+All images were labeled with bounding boxes around the grape bunches and subsequently cropped.
+The resulting images vary in size, as some are close-ups while others contain multiple targets.
+The combination of season, sensor, and gray card presence creates differences summarized in Table 4.
+The rows in the table represent atomic splits—groups with the same data distribution—and the variations between them introduce the covariate shifts we aim to study.
+Table 4: Description of the atomic splits according to the device, image format, gray card presence, and the possibility of having expert ground truth from white-balanced images.
+Dataset Id Year Smartphone Reflex d435i Num images Num samples Expert GT Color GT Type Jpg Raw Card Raw Card RGB IR Card A 2021 motog8 yes no no no - no no - 100 100 no no B 2021 motog8 no yes yes yes yes no no - 38 38 possible yes C 2021 motog8 yes yes yes yes yes no no - 62 62 possible yes D 2021 motog8 yes yes yes yes yes no no - 50 50 possible yes E 2022 realmi yes yes no no - yes no no 1 2 no no F 2022 realmi yes yes yes no - yes no yes 6 22 possible yes G 2022 realmi yes yes no no - yes no yes 1 2 no no H 2022 realmi yes yes yes no - yes no yes 4 19 possible yes I 2022 motog8 yes yes yes no - yes yes yes 8 51 possible yes
 Figure 3: Example of an image collected with a gray card for white balance correction.
 An agronomist performs on-field measurements by randomly sampling the SSC of berries from the top, middle, and lower parts of the bunch and then averaging the values.
-Name Split # samples SSC mean SSC std SSC MAD Description
-$T1_{STD}$ 112 17.19 2.43 1.99 Phone images in JPG format captured in the 2021 season
-$T1_{WB}$ 112 17.19 2.43 1.99 Phone images color corrected from RAW captured in the 2021 season
-$T2_{STD}$ tr 212 16.23 2.52 2.02 Phone images captured in JPG format during the 2021 season.
-$T2_{STD}$ te 92 18.91 2.16 1.71 Phone images captured in JPG format during the 2022 season.
-$T2_{WB}$ tr 150 17.54 2.46 2.04 Phone images captured in RAW format during the 2021 season.
-$T2_{WB}$ te 92 18.91 2.16 1.71 Phone images captured in RAW format during the 2022 season.
-$T2_{ALL}$ tr 300 17.54 2.46 2.04 Phone and reflex images in RAW format, 2021 season.
-$T2_{ALL}$ te 92 18.91 2.16 1.71 Phone and reflex images in RAW format, 2022 season.
-$T3_{STD}$ tr 352 17.67 2.75 2.33 All jpg images, across all seasons and devices.
-$T3_{STD}$ te 44 15.96 1.68 1.36 The test set contains images with SSC content close to the threshold (16° Brix).
-$T3_{WB}$ tr 440 18.27 1.98 1.96 All white balanced images, across all seasons and devices.
-$T3_{WB}$ te 44 15.96 1.36 1.36 The test set contains images with SSC content close to the threshold (16° Brix).
-Table 1: $T1_{STD}$ and $T1_{WB}$ consist respectively in jpg and raw versions of images collected with a smartphone camera during the 2021 season.
-$T2$ datasets are images collected with a smartphone camera, but where the training splits are from the 2021 season, and the test set from the 2022 season.
-In this case, there is no exact correspondence between $T2_{STD}$ and $T2_{WB}$ jpg and raw images.
-The $T3$ datasets serve a different purpose, as they combine images from the $T1$ and $T2$ datasets, mixing data across all seasons and devices, specifically the smartphone, Reflex (only for the WB variant) and d435i cameras.
-The $T3$ training set has more samples than the $T2$.
-At the same time, the $T3$ test set consists of images with SSC threshold quite close to the ripeness level for harvesting (16° for Pizzutello), making it significantly more challenging than the training set.
-Labels: All the images were labeled both for SSC and color.
-For SSC, we used an Atago Hikari PAL2 digital refractometer, capable of measuring berry sugar content (degree Brix, °Brix) by contact without damaging the berry itself.
-This instrument has a range from 9° to 25°.
-Typically, depending on the grape variety, an SSC of 16°-18° is considered good enough for harvesting.
+Figure 4: Some examples of paired JPG and RAW (white-balanced) images collected with the same phone device.
+Labeling protocol: All the images were labeled both for SSC and color.
+For SSC, we used an Atago Hikari PAL2 digital refractometer, capable of measuring berry sugar content in °Brix by contact without damaging the berry itself.
+This instrument has a range from 9° to 26°.
+Typically, depending on the grape variety, an SSC of 16°-18° is considered high enough for harvesting.
 Since we are trying to assess how well the estimation based on appearance would perform compared to human performance, we adopted a labeling procedure similar to the SSC sampling performed by agronomists.
-Figure 2 visually illustrates the process.
+Fig.
+3 visually illustrates the process.
 Every berry in a grape bunch has a different SSC, and the sugar concentration in each berry is also not uniform due to gravity.
@@
 This color index accounts for color saturation and distribution across the berries and the bunch.
-An example of the color variability is given in Fig: 3.
+An example of the color variability is given in Fig.
+5.
 Figure 5: Examples of the six color levels for color classification.
 A clear distinction between categories may be challenging for an untrained eye.
+Agronomist classify the color category taking into account both the general color shades of the bunch and the density of highly colored berries.
 However, our labeling validation tests described in Section 3.1.2 indicate consistent labeling despite some uncertainty in the classification of neighboring classes.
-2.2 Estimation with Histogram Features
-The main use-case for the estimators we design in this work is to be deployed on a robotic platform, embedded edge devices, or smartphone-like device, thus with limited memory and computational power.
+Experimental Dataset Splits: Our dataset is built upon a foundation of atomic splits, which are minimal, homogeneous groups of images categorized by season, device, and processing type, as detailed in Table 4.
+We combined these atomic splits to create several experimental datasets designed to evaluate model performance in specific scenarios.
+These combined datasets are distinguished by their image type: standard JPG images (indicated by the STD subscript) or white-balanced images derived from raw captures (indicated by the WB subscript).
+A summary of these experimental datasets is presented in Table 5.
+The datasets are grouped by their analytical purpose:
+  • 1.
+    The Simple splits are small, single-season (2021) datasets used for initial, exploratory assessments of our models. They do not have a predefined train/test division.
+  • 2.
+    The Season21ToSeason22Phone splits fix the sensor type to smartphones and are designed to evaluate cross-seasonal performance by training on data from one season and testing on the next.
+  • 3.
+    The MixedSeason splits represent the most challenging scenario.
+They are designed to simulate a real-world deployment by training on smartphone and Reflex images (mimicking data collection by agronomists) and testing on d435i images (mimicking a robotic platform).
+This directly tests the models’ robustness to the covariate shift induced by different acquisition devices.
+To provide a quantitative baseline for our experiments, Table 6 summarizes key statistical properties of these datasets.
+The table details the size of each train and test split, along with the natural variability of the SSC labels, measured by the mean, standard deviation, and Mean Absolute Deviation (MAD).
+These statistics will serve as a reference for evaluating the performance of our proposed estimation models.
+Table 5: Summary of the experimental dataset splits used for model training and evaluation.
+[…]
+Table 6: Statistical properties of the experimental dataset splits.
+[…]
+2.2 SSC Estimation Models
+In this Section, the two algorithmic approaches to SSC estimation are described.
+2.2.1 Estimation with Histogram Features
+The main use-case for the estimators we design in this work is to be deployed on a robotic platform, embedded edge devices, or smartphone-like devices, thus with limited memory and computational power.
 For this reason, the first approach is based on classic hand-crafted features, specifically on histograms.
@@
 A key observation that can be made with this setup is that even with tight bounding boxes from the detector, some regions will often contain background information due to the shape of grape bunches.
-To address this issue, we modified the feature processing to retain only specific regions in a ”cross-shaped” selection, as illustrated in Figure 4.
-This selection focuses on the areas of the image where grape content is most likely to be present.
-Figure 4: Each grape bunch in the scene is cropped based on the bounding box provided by the detector.
-The resulting image is then binned, with some bins dropped as they contain background information with high probability.
-This picture shows an example, but the actual bin size and dropping schemes are hyper-parameters described in Section 2.
-In this simple example, the n-uple defining column and row ranges described by Eq.1 would be [1,3,1,2]
-To simplify hyper-parameter exploration, we defined three approaches for region selection: a ”fat” cross-shaped pattern, a ”thin” one, and one using all regions without any exclusion.
-More in detail, the three configurations are defined by the starting and ending cell in the column and row direction as follows:
+Adding a background subtraction deep network, such as in Qin et al., [14], could have been effective to some extent, at the cost of changing the scale of the computational costs.
+For this reason, we decided to opt for a more straightforward approach.
+To address this, we developed a simple heuristic to remove regions likely to contain background pixels.
+Based on the empirical observation that grape bunches are typically wider in the middle and narrower at the top and bottom, we retain only a ”cross-shaped” selection of regions, as illustrated in Fig.
+6.
+We treat the exact dimensions of this cross-shaped pattern as a hyper-parameter to be optimized.
+To simplify this exploration, we defined three configurations in Equation 1: none (keeping all regions), fat (dropping corner regions), and thin (dropping corners and more bottom regions).
+More in detail, the three configurations are automatically defined by the selection of the number of horizontal ($n_{bin_{x}}$) and vertical ($n_{bin_{y}}$) divisions (the width and height of the grid in Fig.
+6) as follows:
 […]
@@
 […]
-We compared these approaches to determine the most effective feature selection strategy. In summary, the hyper-parameters that were considered while designing the features are:
-  • 1.
-    Image starting size (width, height)
-  • 2.
-    Number of cells in both the width and height directions
-  • 3.
-    The cross pattern used to remove the background
-  • 4.
-    Use of HSV instead of RGB channels
-We performed a hyper-parameter search for both the task of SSC and color degree estimation using Ridge Regression.
-For brevity, since color estimation is an easier task to perform, in Section 3 we omit the details of the experiments for color estimation, while the SSC ones are explored in depth.
-2.3 Estimation with Multi-Task Learning DNN Architecture
-As mentioned before, one of the research questions we are exploring is the correlation between the color of the berries and sugar content.
-The correlating latent factors are many and difficult to model explicitly.
-Among them are climate, weather, sunlight exposure, pruning strategies, and the use of nitrogen and other chemicals.
-Given this complex interaction between different factors, we decided to design a neural architecture that could extract common features to be used for both estimation tasks.
-Our model is shown in Figure 5.
+Averaging the RGB values on each cell and concatenating them produces a final feature vector with length of $3\times n_{bin_{x}}\times n_{bin_{y}}$.
+In our experimental validation, We considered also the exploration of the effects of the initial image size and the use of Hue-Saturation-Value (HSV) color space.
+Given the small number of images, to avoid overfitting, we also considered $L2$ regularization.
+We use Ridge Regression as the estimation algorithm, since it is the simplest interpretable regression algorithm that allows for regularization.
+Clearly, other more sophisticated models – such as Support Vector Machines of Random Forests – can be used for better estimation results, but in this work we are interested in exploring the characteristics of the features and not only the peak performances.
+Figure 6: An illustration of the ”cross-shaped” feature selection process for histogram-based estimation.
+An input image is divided into a grid of $n_{bin_{x}}$ by $n_{bin_{y}}$ bins ($5\times 5$ grid shown).
+To heuristically remove background pixels, two patterns are tested: a fat pattern that retains a wider central cross-section and a thin pattern that is narrower at the bottom, mimicking the typical shape of a grape bunch.
+The indices of the kept bins are determined empirically using Equation 1.
+For the $5\times 5$ example shown, this results in the index ranges [1,3,1,3] for the fat pattern and [1,3,1,2] for the thin pattern.
+2.2.2 Estimation with Multi-Task Learning DNN Architecture
+The SSC and color distribution can be viewed as conditioned on the environmental and agronomic parameters of the field we are considering.
+While both the color and the SSC depend on the internal chemical state of the berries, this state itself is influenced by macroscopic environmental and managerial factors such as growing conditions, training practices, and nutrient availability.
+The overall appearance of the grape bunches is also affected by these factors, which we assume largely homogeneous with respect to key management practices and environmental conditions.
+Nevertheless, our proposed methods are designed to be robust to the inherent noise and micro-variations arising from individual plant differences, which are expected in any real-world field data.
+The architecture of the multi-task network, illustrated in Fig.
+7, reflects this reasoning.
 To distill a feature vector, we employed an Auto-Encoder (AE) architecture, instead of using only a feature extraction backbone alone.
+The encoder component captures common features in appearance variations that may correlate with the internal chemical state of the grapes.
+These shared features are then leveraged by two separate prediction heads to estimate the color class and SSC in parallel.
 The reason is that we are in a data-scarcity scenario, so having the additional reconstruction loss of the auto-encoder, focusing on appearance, helps in strengthening the back-propagation signal and reduces over-fitting.
-Our auto-encoder is a UNet with a ResNet18 backbone.
+Our auto-encoder is a UNet [15] with a ResNet18 [7] backbone.
+The two heads share a common ”neck,” a CNN with 512 output channels, ensuring part of the representation is shared between the tasks.
+Each head has a symmetrical structure, comprising three CNN blocks that progressively reduce the feature size, followed by adaptive average pooling Zhou et al., [27] and a 1x1 convolution Lin et al., [11] for the final estimation.
 In general, UNet-like architectures alone are not suitable for reconstruction problems, since the skip connections would cause the network to simply learn an identity function.
@@
 The two heads have the same structure, but in the case of color estimation, the outputs are logits for the multi-class problem, while for the SSC estimation are the actual estimations.
-The loss for the architecture is the combination of the reconstruction loss of the AE (pixel-wise Mean Squared Error (MSE)), with the addition of a Mean Absolute Error (MAE) loss for the SSC estimate, and a Cross-Entropy loss for the color classes:
+Figure 7: The multi-task network architecture used for SSC and color estimation.
+The model is based on a U-Net (Ronneberger et al., [15]) with a ResNet18 (He et al., [7]) backbone, which acts as an auto-encoder to learn a compact image representation (outputting the reconstruction loss, $\mathcal{L}_{rec}$).
+Features from the Bottleneck Conv layer are passed to a block of Shared features.
+From here, the network splits into two parallel heads: a Brix regressor for SSC estimation (outputting $\mathcal{L}_{brix}$) and a Color classifier for quality grading (outputting $\mathcal{L}_{color}$).
+(Diagram generated with PlotNeuralNet [8]).
+The loss for the architecture is a classical compound loss that combines the reconstruction loss of the AE (pixel-wise MSE), with the addition of a Mean Absolute Error (MAE) loss for the SSC estimate, and a Cross-Entropy loss for the color classes:
 $\mathcal{L}_{MT}=\mathcal{L}_{rec}+\alpha\mathcal{L}_{color}+\beta\mathcal{L}_{brix}$ (2)
 where $\alpha$ and $\beta$ are weighting factors to balance the different contributions of the three losses, whose values have been cross-validated together with other hyper-parameters.
-Figure 5: The multi-task network is based on a U-Net [13] with a ResNet18 [7] backbone, designed to encode the image into a compact representation.
-Features extracted from the bottleneck layer are fed into a two-head estimator: one for SSC (regression) and another for color (classification).
-The two heads share a common ”neck,” a CNN with 512 output channels, ensuring part of the representation is shared between the tasks.
-Each head has a symmetrical structure, comprising three CNN blocks that progressively reduce the feature size, followed by adaptive average pooling [24] and a 1x1 convolution [10] for the final estimation.
-2.3.1 Hyper-parameter Validation
+To optimize the hyper-parameters of our multi-task DNN, we developed a custom scoring metric designed to balance statistical accuracy with agronomic needs.
+We call this the Brix-Color-Hinge (BCH) metric, as it integrates the regression error for °Brix, the classification accuracy for color, and includes hinge-like penalties for predictions that fall below minimum quality thresholds.
+This approach ensures that the selected model is not only accurate but also practical for harvesting decisions.
+The detailed mathematical formulation of the BCH metric is provided in C.
+3 Experiments
+This Section describes the experiments of the two proposed algorithmic solutions to SSC estimation.
+In order to give perspective to these results, some experiments were performed to set a human performance baseline.
+For this reason, the Section starts with the explanation on how human performances were tested in this scenario (Section 3.1), then proceeds with the actual experiments of the histogram feature model (Section 3.2) and multi-task DNN model (3.3).
+Both algorithmic solutions have been initially developed and trained locally on a laptop equipped with a AMD Ryzen 9 5980HX Mobile CPU and a NVIDIA GeForce RTX 3080 GPU with 16GB of VRAM, running Ubuntu 22.04, while the hyper-parameter optimization for all models have been run on a NVIDIA DGX workstation equipped with A100 GPUs.
+All the code has been written in python3.
+The data and code for all the experiments can be found on the accompanying GitHub repository [^1] https://github.com/Smart-Agri-DIAG/robots-tasting-grapes.
+3.1 Human Performance Baselines
+3.1.1 Human Evaluated SSC and Color Correlation
+To establish an initial human performance baseline and assess experts’ ability to estimate ripeness from appearance, we performed a specific test.
+We presented $100$ grape images to a number of agronomists with experience in table grape ripeness estimation.
+For these images, a °Brix ground truth was collected with the technique described in Fig.
+3.
+We used the JPG smartphone images collected during the first data collection session (24 September 2021) for Pizzutello Nero.
+The images were presented without further context, and the experts were asked to assign a numerical value, on a scale from 1-to-10, to the degree of color, peduncle lignification stage, and berry shape, respectively indicated in the following as color_score, lignification, and berry_shape.
+Note that this 1-to-10 color scale differs from the 1-to-6 scale defined in Section 2.1.2; This experiment utilized a decimal scale for ease of human interpretation.
+To create a binary harvesting prediction from this data, we used the agronomists’ color_score as a proxy for their final decision, applying a threshold of 6.
+We called the thresholded prediction harvesting_pred.
+We then measured the correlation between all three visual estimates and the measured mean_brix to validate our core hypothesis.
+The correlation scores are shown in Fig.
+8.
+The correlation score between the average °Brix (mean_brix) and color evaluation (color_score) is $0.505$, with $p<0.001$ and $CI=[0.342,0.638]$, the correlation between mean_brix and lignification (lignification) is $0.337$, $p<0.001$ and $CI=[0.151,0.501]$, while the correlation with berry shape (berry_shape) score is considerably lower, at $-0.044$, which is below any significance threshold.
+The observed correlation between human expert visual color assessment and actual measured °Brix values suggests that a data-driven algorithm could achieve human-level accuracy in estimating grape quality for harvesting decisions.
+Figure 8: Correlation matrix of key parameters from the human evaluation study on 100 grape images.
+The matrix shows the relationship between the measured mean_brix and the scores assigned by experts for color_score, lignification, and berry_shape.
+The strong correlation of 0.505 between mean_brix and color_score is the key finding that validates using visual appearance for ripeness estimation.
+This experiments also helps in setting a human performance baseline.
+Given that a °Brix value greater than $17$ is considered sufficient for harvesting, we use the thresholded mean_brix as the harvesting ground truth label, comparing it to harvesting_pred.
+The precision and recall values were $0.43$ and $0.83$, respectively, translating to a false discovery rate (Type I error rate) of $0.57$ and a false negative rate (Type II error rate) of $0.17$.
+These values will be compared with similar scores computed with the proposed algorithms in the following Sections.
+As a final note to this analysis, it is important to note that in the case of harvesting, Type I errors (false positives) carry a higher cost than Type II errors.
+This is because ripe bunches left on the vine will eventually be harvested, whereas unripe harvested bunches must be discarded.
+By collecting table grape bunches with a color class above the threshold of $6$ (the definition for the harvesting_pred attribute) a significant number of unripe bunches are collected due to a mismatch between °Brix and color.
+This raises the question of whether increasing the threshold for harvesting_pred could trade Type I errors for Type II errors.
+However, computing the precision and recall scores for a harvesting_pred value obtained with a threshold of $7$ yields $0.35$ and $0.37$, respectively, indicating an increase in both error types.
+This test shows that, while °Brix and color are correlated across the broad range of the color scale used in this experiment, at the same time, the relationship is not strictly linear.
+There are a considerable number of both insufficiently colored bunches with sufficient °Brix and highly colored bunches with insufficient °Brix.
+The high number of bunches in a borderline situation that the experts miss-classify shows that training an appearance-based estimator is a non-trivial task.
+3.1.2 Color Constancy effect on Human Evaluation of Color
+As part of our exploratory analysis, we investigated whether the camera’s automatic post-processing (standard JPG) versus a color-corrected image (white-balanced RAW) would significantly affect a human expert’s color evaluation.
+To test this, we conducted a study where agronomists labeled 50 pairs of images (one standard, one white-balanced) in a randomized sequence.
+A statistical analysis (z-test) of the assigned labels revealed no significant difference between the two groups (p = 0.17).
+This null finding confirmed that for human experts, the two image types were perceptually similar and justified our subsequent approach of testing our computer vision models on both standard and white-balanced images to evaluate their robustness.
+3.2 Histogram-based Model Evaluation
+To establish a performance baseline for a computationally inexpensive approach, we evaluated the histogram-based feature model using Ridge Regression.
+The experiments were conducted on the Simple dataset splits, which contain 112 samples.
+For this analysis, the data was divided into a training set (89 samples) and a final test set (23 samples).
+The training images were augmented with horizontal flips, resulting in a total of 178 training images.
+As previously mentioned, both standard JPG and white-balanced images were evaluated to assess the influence of post-processing on the feature-based algorithms.
+Model performance and hyper-parameters were then assessed using a 5-fold cross-validation procedure on the training set.
+The model’s performance was optimized by conducting a grid search over its key hyper-parameters.
+These included the number of horizontal and vertical histogram bins ($n_{bin_{x}}$, $n_{bin_{y}}$), the resulting feature vector length (feature_len), the choice of color space (RGB, HSV, or CIE-Lab), and the regularization strength ($\lambda$).
+The full list of parameters and their search ranges is detailed in A.
+The best-performing model achieved a validation MAE of $1.46$Brix.
+This result represents a notable $\mathbf{27\%}$ improvement over the statistical baseline for this dataset, which had a Mean Absolute Deviation (MAD) of $1.99$Brix.
+The analysis of the hyper-parameter search revealed key insights into the model’s behavior.
+As shown in the feature correlation plot (Fig.
+9(a)), the regularization strength ($\lambda$) and parameters affecting the feature_len had the most significant impact on performance.
+The analysis also highlighted the consistent benefit of using an alternative color space, with both HSV and CIE-Lab generally outperforming standard RGB.
+Furthermore, we investigated the effect of using standard JPG images versus color-corrected, white-balanced (WB) images.
+Fig.
+9(b) shows that models trained on the standard JPG images consistently achieved a lower validation error than their counterparts trained on white-balanced images with identical hyper-parameters.
+This suggests that for this feature-based approach, the camera’s default processing pipeline may contain beneficial information, and the extra step of color correction is not necessary.
+Figure 9: a) Correlation between validation error (MAE) and hyper-parameters for the top 10% performing runs.
+This plot highlights the direction (positive or negative) of the correlation.
+Parameters controlling model complexity, such as feature_len and lambda (regularization strength), have the strongest correlation with performance.
+b) Test error for models with hyper-parameters identical to the best model, but with varying lambda values (plotted on a logarithmic scale) and color correction.
+This plot isolates the effect of regularization and demonstrates that, for this model, standard JPG images consistently yield better performance than their color-corrected counterparts.
+3.3 Multi-task DNN Model Evaluation
+[…]
+An important observation from the multi-task training process is that the optimal performance for SSC estimation (regression) and color quality (classification) were not always reached at the same training epoch.
+This highlights an inherent trade-off when using a single feature extractor for multiple correlated but distinct tasks.
+This also implies that, for a specific application where the top performance of either SSC or color estimation is required, one could select different model checkpoints depending on whether the primary goal is maximizing SSC accuracy or color classification accuracy.
+Finally, we analyzed the impact of color correction on the DNN model’s performance, similarly to what was done with the histogram features.
+To isolate this effect, we selected the best-performing models on the standard $\texttt{MixedSeason}_{\texttt{STD}}$ dataset and compared them against runs using identical hyper-parameters but trained on the white-balanced $\texttt{MixedSeason}_{\texttt{WB}}$ data.
+Fig.
+10 plots the results of this comparison against the log-ratio of the regularization weights ($\alpha/\beta$).
+The resulting regression lines show that the performance difference between the two sets of images was minimal, with largely overlapping confidence intervals.
+This suggests that the deep learning model is robust to these color variations and can learn the relevant features effectively without requiring complex color calibration steps prior to deployment.
+Figure 10: Linear regression analysis of runs with hyper-parameters similar to the top 20 runs on the $\texttt{MixedSeason}_{\texttt{STD}}$ and $\texttt{MixedSeason}_{\texttt{WB}}$ datasets.
+The shaded area represents the $95\%$ confidence interval.
+While the models trained on color-corrected samples present slightly higher values on average, the smaller difference compared to Fig.
+9(b) suggests no practical advantage in using standard images over the white-balanced ones in this context.
+4 Discussion
+The results of our experiments demonstrate that low-cost, in-field RGB imaging can be a viable alternative to more complex and expensive systems for SSC estimation.
+In this section, we synthesize the performance of our models, compare them against the established baselines and the state-of-the-art, and discuss the practical implications and limitations of our work.
+4.1 Summary of Performance and Comparison
+To provide a holistic view of our findings, Table 7 consolidates the performance of our proposed models against the statistical and human baselines, as well as against other non-contact methods in the literature.
+Table 7: Master comparison of model performance against baselines and state-of-the-art methods.
+[…]
+Our deep learning model clearly learned a meaningful pattern from the images, achieving a best MAE of 1.55 °Brix on the Season21ToSeason22Phone dataset and 1.05 °Brix on the challenging cross-device MixedSeason dataset.
+This significantly outperforms the naive statistical baselines of 1.71 and 1.36 °Brix MAD respectively for that same test set.
+To better address how these errors translate to harvesting process performance and to benchmark against practical expertise, we also evaluated our DNN model on the binary harvesting task.
+As shown in Table 7, our model achieved a precision/recall of $\mathbf{0.91}/\mathbf{0.84}$ on Season21ToSeason22Phone split and of $\mathbf{0.62}/\mathbf{0.83}$ on the MixedSeason split.
+This is compared to the experts’ performance of $0.43/0.83$.
+Being able to keep a high recall, while showing a higher precision is particularly valuable in a real-world scenario, as it would reduce the costly error of harvesting unripe grapes.
+In addition, our experiments consistently showed that the F1 score computed from these Precision and Recall values is highly correlated to the BCH metric (in most runs the maxima of the F1 score correspond to the minima of the BCH metric) suggesting that the model’s high performance derives from the effective combination of features related to both color and SSC estimation.
+When benchmarked against other visual estimation techniques, our low-cost RGB approach demonstrates highly competitive performance.
+While direct comparison is challenging due to different grape varieties and conditions, our best MAE of 1.05 °Brix is favorable when compared to reported RMSE values from in-field HSI systems, which range from 1.27 to 2.20 °Brix.
+It also represents a significant improvement over other in-field RGB methods that report errors closer to 4.63 °Brix RMSE, and approaches the accuracy of methods performed under controlled laboratory conditions.
+4.2 Practical Implications and Trade-offs
+Our work presents two distinct solutions tailored for different operational needs.
+The histogram-based model, while less accurate, offers significant advantages in computational efficiency and simplicity.
+Once trained, the estimation reduces to a dot product between the feature vector and the estimator weights.
+The most expensive operation in that case is the pixel-wise integral sum needed for the histogram computation, which has a computational complexity of O(N), where N is the number of pixels.
+However, as it was shown in the experiments, given the low frequency content of the grape image it is possible to work with very low resolution images and drastically reduce N.
+To summarize, the histogram features’ low memory and processing footprint make it suitable for real-time deployment on resource-constrained edge devices or as a first-pass filter on a robotic platform.
+In contrast, the DNN model provides superior accuracy and generalization, particularly across different devices and seasons, at a greater computational cost.
+This makes it ideal for applications where performance is paramount, such as a centralized quality control system or deployment on robots with dedicated GPU hardware.
+4.3 Limitations and Future Work
+We acknowledge several limitations that provide avenues for future research.
+This study was conducted on a single black grape variety, Pizzutello Nero.
+While this is common in literature (see Table 7), to extend it to a white variety some adjustments in the setup are required.
+In the case of black or red grape varieties, the color gradient can be perceived as a contrast change, while in with the white varieties it is a hue and saturation change.
+To detect variations of the latter kind, a color calibrated RGB imaging system is needed, or, in alternative, to have a calibration pattern in the scene.
+This is not impossible to achieve and it is an interesting future development for this work.
+Additionally, while our methods showed robustness to the uncontrolled lighting of the field, further experiments under more extreme weather conditions (e.g., direct bright sunlight vs.
+heavy clouds) could better quantify the operational domain of the system.
+Future work will focus on expanding the dataset to include these variations and on exploring multi-sensor fusion techniques to further improve robustness.
+5 Conclusions
+This work successfully demonstrated the feasibility of estimating Soluble Solid Content (SSC) in grapes using low-cost RGB sensors in challenging, uncontrolled field conditions.
+By creating a novel multi-season, multi-device dataset of Pizzutello Nero labeled for both SSC and color, we established a framework for developing and validating data-driven estimation models, addressing the practical limitations of expensive laboratory or hyperspectral equipment.
+While numerous studies explore laboratory equipment and hyperspectral data for SSC estimation through visual sensors, this study is among the few to focus exclusively on the practical scenario of RGB data collected directly in field conditions.
+The complexity of this task arises from covariate shifts inherent to the problem, including seasonal variations and differences in imaging devices.
+Two algorithmic approaches were proposed: a lightweight method using classical handcrafted features, suitable for devices with limited computational power, and a multi-task deep neural network architecture that, while more computationally demanding, effectively leverages correlations between visual appearance (color and shades) and SSC.
+Both methods were experimentally validated using field-collected data.
+Results indicate that the proposed techniques achieve reliable SSC estimation with a reasonable margin compared to the average human estimator.
+Our multi-task deep neural network achieved a best-performing Mean Absolute Error (MAE) as low as 1.05 °Brix on a challenging cross-device test set.
+This level of accuracy is highly competitive with reported results from in-field hyperspectral systems and is comparable to our human expert baseline, whose harvesting decisions the model could match with high precision and recall.
+The demonstrated robustness to seasonal and hardware-induced covariate shifts position our solutions as viable prototypes for low-cost, on-field applications.
+To tackle this challenge, a dataset spanning multiple years was collected, incorporating data from various camera devices and annotated for SSC and color quality based on agronomic practices.
+Human performance in assessing the correlation between color and SSC was evaluated through two user studies, which informed the design of data-driven algorithms for estimating both SSC and color.
+To foster further research we release our complete dataset, code, trained models, and training scripts used in this study.
+Future work will focus on extending the approaches to green cultivars, and on further studying how low-cost estimators can directly approximate hyperspectral devices in agronomic contexts..
+CRediT authorship contribution statement
+Thomas A.
+Ciarfuglia Conceptualization, Methodology, Software, Validation, Formal analysis, Data Curation, Investigation, Resources, Writing - Original Draft, Writing - Review & Editing, Visualization, Supervision, Project administration, Funding acquisition.
+Ionut M.
+Motoi: Software, Validation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization Leonardo Saraceni: Software, Validation, Data Curation Daniele Nardi: Funding acquisition, Writing - Review & Editing
+Declaration of competing interest
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Acknowledgments
+[…]
+This work is part of a project that has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 101016906 – Project CANOPIES
+[…]
+This work has been partially supported by project AGRITECH Spoke 9 - Codice progetto MUR: AGRITECH ”National Research Centre for Agricultural Technologies” - CUP CN00000022, of the National Recovery and Resilience Plan (PNRR) financed by the European Union ”Next Generation EU”.
+This work has been partially supported by Sapienza University of Rome as part of the work for project H&M: Hyperspectral and Multispectral Fruit Sugar Content Estimation for Robot Harvesting Operations in Difficult Environments, Del.
+SA n.36/2022.
+Appendix A Hyper-parameters Exploration for the
+Histogram-based Estimators
+Following from Section 3.2, to allow full reproducibility of the experiments, here we give the details of the hyper-parameter search for the histogram features.
+Table 8 summarizes the hyper-parameter ranges investigated in this experiment.
+Table 8: Hyper-parameters ranges for the histogram feature-based models.
+Hyper-parameter Range description
+$n_{bin_{x}}$ 4, 8, 16 number of image columns bins.
+$n_{bin_{y}}$ 8, 16, 32 number of image rows bins.
+$cross$ ”none”, ”thin” or ”fat” How many bins to drop for background elimination. See Eq.1 for details.
+$color\_space$ ”none”, ”hsv” or ”Lab” Use of CIE-La*b*, HSV or RGB color coordinates.
+$dataset$ low_res, med_res, high_res, Image resolution, corresponding to $320\times 240$, $640\times 480$ and $1280\times 720$ respectively, in color corrected and standard form.
+wb_low_res, wb_med_res, wb_high_res
+$\lambda$ 0.33, 1, 3, 9, 27, 81, 243, 729 Ridge Regression regularization parameter.
+Given the manageable number of combinations and the low computational cost of each experiments, we performed a full grid search to find the optimal hyper-parameters.
+To identify the most relevant ones we used the importance factor, as described in Grömping, [5] and implemented in the Weights&Biases software suite for experiments tracking [3].
+Fig.
+11 gives an example of this analysis, while Table 9 shows some of the best performing runs.
+The full results and hyper-parameters for all 3888 experimental runs are publicly available for inspection to ensure full transparency in the GitHub11footnotemark: 1 and Weights&Biases [^5] https://wandb.ai/sparviero19/CQEM-193_Brix_sweep/reports/CQEM-193-Sweep-results-on-sweep-zgyffmc5--Vmlldzo2Njc3MjA4 companion sites.
+Table 9: Top 10 histogram-based models, ranked by validation mean absolute error (Val MAE), along with their corresponding hyper-parameters values.
+Test MAE Val MAE Train MAE $\lambda$ $n_{bin_{x}}$ $n_{bin_{y}}$ $cross$ $color\_space$
+1.46 2.10 1.58 9 16 8 none HSV
+1.51 2.13 1.55 9 8 16 none HSV
+1.52 2.14 1.61 3 8 8 none HSV
+1.51 2.14 1.18 9 8 32 none HSV
+1.52 2.16 1.62 1 16 8 thin CIE-Lab
+1.49 2.16 1.25 3 16 8 none HSV
+1.52 2.18 1.43 3 16 8 fat HSV
+1.52 2.19 1.72 0.33 8 8 thin CIE-Lab
+1.43 2.20 1.19 3 8 16 none HSV
+1.49 2.33 1.22 1 16 8 thin HSV
+Figure 11: Training a Random Forest using the hyper-parameter values to predict the validation error and then extracting the importance scores is a common way to extract the relative impact of each feature on the final estimator performance.
+Here we can see that the choice of initial image resolution is not relevant, while the regularization, the number of bins (that influence feature length and flexibility), and the color space have to be carefully selected.
+Table 10: Hyper-parameter descriptions for the Multi-task experiments.
+Hyper-parameter name Range description
+$\alpha$ and $\beta$ 1, 10, 50, 100, 500, 1000 The loss balancing factors (Eq. 2).
+$batch\_size$ 32, 64, 128 Training batch size.
+$learning\_rete$ 0.00001, 0.00005, 0.000001 Initial learning rate, to be modified by the scheduling strategy.
+$momentum$ 0.8, 0.9, 0.95, 0.99 Momentum for the SGD optimizer.
+$im\_height$ and $im\_width$ 64, 128, 256 The image starting sizes before feature extraction by the auto-encoder.
+$blur\_\sigma_{min}$ and $blur\_\sigma_{max}$ $\left[0.01,0.1,0.5\right]$ and $\left[2.0,4.0\right]$ Together they define the range for the Gaussian blur augmentation that is always applied.
+$color\_space$ ”none”, ”hsv” or ”Lab” Use of CIE-La*b*, HSV or RGB color coordinates.
+$one\_cycle$ $True$ or $False$ Whether to use of not the one cycle scheduling strategy (Smith and Topin, [16]).
+$pretrained$ $True$ or $False$ Whether the resnet18 backbone is pretrained on ImageNet (He et al., [7]) or not.
+[…]
+Appendix B Hyper-parameter Exploration for the Multi-task DNN
+This appendix follows the same principles described in the previous one.
+Table 10 gives a summary of the hyper-parameter ranges explored in the sweep.
+A wide range of hyper-parameters was explored using random search.
+In this case, the dataset splits used are not to be considered proper hyper-parameters, but as a defining subset of the parameter sweep, since the distribution shifts between them are considerable.
+They are included in this table for ease of reference.
+Given the huge parameter space, this time the sweep was performed with random sampling, with a total run count greater than 4000.
+To understand which parameters were most influential, we analyzed the top 25% of all experimental runs.
+An example of the correlation between some of the hyper-parameters and the BCH validation score for this subset is visualized in Fig.
+12.
+The full analysis revealed that standard deep learning parameters like learning rate and regularization weights, along with image dimensions, were the most influential factors on performance.
+The detailed hyper-parameter configurations for the best models are available in our GitHub repository and the complete list of experimental run can be found on the Weights&Biases report.222 https://api.wandb.ai/links/sparviero19/s2p10p3a
+[…]
+Appendix C Multi-Task DNN Hyper-parameter Validation Metric
 Hyperparameter selection in multi-task learning requires identifying a single metric $\mathcal{S}$ that effectively summarizes performance across different models, which is particularly challenging when tasks differ significantly.
@@
 The final score, $\mathcal{S}$ integrates the normalized performance metrics and the agronomic quality criteria.
+We call this metric Brix-Color-Hinge (BCH) metric.
 The formal definition of the score $\mathcal{S}$ is the following:
@@
 $t_{brix}$ and $t_{color}$ represent the agronomic quality thresholds for °Brix and color, respectively, and $\eta$ is the normalization factor between SSC and color scores.
-In our experiments, $t_{brix}=2.5$Brix, which is comparable with the standard deviation of the training data, and $t_{color}=0.8$, since an error rate of $80\%$ is usually considered acceptable for human performance in this context [14].
+In our experiments, $t_{brix}=2.5$Brix, which is comparable with the standard deviation of the training data, and $t_{color}=0.8$, since an error rate of $80\%$ is usually considered acceptable for human performance in this context The Canopies Consortium, [17].
 The normalization factor is empirically set as
 $\eta=\frac{t_{brix}}{t_{color}}.$ (6)
-3 Experiments
-In designing the experimental setup, we had two main aims.
-The first one was to give an initial answer on factors are relevant in estimating SSC from appearance, in order to better understand the problem for future work and data collection campaigns.
-The second aim is to validate the proposed features and estimator models.
-This section begins with the validation of the collected dataset, followed by a description of the exploratory experiments.
-Finally, we describe the validation of our proposed features and models, together with the ablation studies.
-The data and code for all the experiments can be found on the accompanying GitHub repository [^1] To be disclosed upon publication.
-3.1 Exploratory User Studies
-3.1.1 Human Evaluated SSC and Color Correlation
-Feature description
-v1, v2, v3 Upper, middle and lower °Brix measurements.
-m_brix Mean °Brix computed from v1, v2, v3.
-m_b_t Rounded m_brix (unused).
-b_class Binary °Brix class for harvesting (Sufficient, Insufficient). Computed from m_brix with threshold $17$°.
-c_val Expert given score for color development after varaison, on a scale from 1 to 10.
-c_class Binary color class for harvesting (Sufficient, Insufficient). Computed from c_val with threshold $6$.
-l_val, r_val Expert given scores for lignification and berry shape development, respectively, on a scale from 1 to 10.
-Table 2: These are the parameters considered by the agronomist in the exploratory experiment to assess the correlation between °Brix and human color evaluation.
-In order to have an initial human performance baseline and to assess the experts’ actual capacity to estimate ripeness from appearance, we performed a specific test.
-We gave to a number of agronomists with experience in table grape ripeness estimation, $100$ grape images, for which a °Brix ground truth was collected with the technique described in Fig.
-2.
-We used the JPG smartphone images collected during the first data collection session (24 September 2021) for Pizzutello Nero.
-The images were given without any other hint to the experts, asking them to express a numeric evaluation, on a scale from 1 to 10, of the degree of color, peduncle lignification stage, and berry shape.
-Note that this color scale is not the same as the one defined in Section 2.1.2.
-The reason is that this experiment uses a human-friendly decimal scale, while the one defined previously is to be used by a classification algorithm.
-In addition to the aforementioned parameters, the agronomists were asked to guess if the °Brix for the sample was high enough for harvesting purposes (in this case, $\geq 17$).
-Table 2 summarizes the experiment variables name and meaning.
-First, we checked the correlation between human-given values of c_val, l_val and r_val and the measured mean °Brix value, m_brix to explore the hypothesis where color and other visual elements could be a good predictor for sugar content.
-In Fig.
-6 the correlation matrix is shown.
-The correlation score between the average °Brix (m_brix) and color evaluation (c_val) is $0.505$, with $p<0.001$ and $CI=[0.342,0.638]$, the correlation between m_brix and lignification (l_val) is $0.337$, $p<0.001$ and $CI=[0.151,0.501]$, while the correlation with berry shape (r_val) score is much lower, $-0.038$ way below any significance threshold.
-The observed correlation between human expert visual assessment and measured °Brix values suggests that a data-driven algorithm could achieve human-level accuracy in estimating grape quality for harvesting decisions.
-Even if there is a significant correlation with the lignification of the peduncle, in this work we focused on the color, leaving the use of peduncle lignification as a predictor for future work.
-The second outcome of this experiment was to define some human harvesting performance baselines.
-For this question, we consider the c_class, the color category given by the agronomists to each sample, as the harvesting prediction and the thresholded values b_class as the binary labels.
-The precision and recall values are respectively $0.43$ and $0.83$, which means a false discovery rate (type I error rate) of $0.56$ and false negative rate (type II error rate) of $0.17$.
-It is important to state that in the case of harvesting, type I errors (false positives) have a higher cost than type II since all the ripe bunches left on the vine tree will be eventually harvested, while the harvested unripe bunches have to be disposed of.
-We see that by simply collecting the table grape bunches with a color class over the threshold of $6$ (the definition for the c_class attribute) a lot of unripe bunches are collected due to °Brix to color mismatch.
-One question that arises then is if by increasing the threshold for c_class, it would be possible to trade type I errors for type II errors.
-Computing the precision and recall scores for a c_class value obtained with a threshold of $7$ for example, we get respectively $0.35$ precision and $0.37$ recall, which is an increase in both error types.
-This test shows that, while °Brix and color are correlated on the broad range of the color scale of this experiment, at the same time, the relationship is not quite linear, i.e.
-there are both a considerable number of insufficiently colored bunches that have sufficient °Brix, and highly colored bunches with insufficient °Brix.
-The high number of bunches on a borderline situation that the experts miss-classify shows that the question of training an appearance-based estimator is a non-trivial one.
-Figure 6: Correlation matrix of the parameters estimated by human experts looking to a sample of 50 images of table grape bunches, as described in Section 3.1.
-The key factor for our evaluation is the correlation between the SSC guess (m_brix) and the evaluation of bunch color (c_val).
-3.1.2 Color Constancy effect on Human Evaluation of Color
-One of the early questions that arose from data collection was to what extent the camera image processing pipeline could influence the possibility of estimating the correct color stage, and therefore influence the °Brix estimate.
-Smartphone cameras, like most camera devices, do not process the raw signal from their CCD/CMOS sensors for radiation accuracy, but more for aesthetic purposes.
-This means that color accuracy is often lost in the nonlinear pipeline that transforms raw images into JPGs.
-A second consideration was related to the difference between the human vision system and a camera device.
-We know that human experts and workers alike are able to guess ripeness from appearance, but to what extent some of the peculiarities of the human vision system influence this guess (e.g.
-consider the brightness constancy effect) is unknown.
-While color perception is a complex topic, we focused our analysis on a specific question: whether human experts assign significantly different color labels when viewing white-balanced versus standard images.
-We selected 50 images from a single field trip that had paired raw and JPG image formats.
-The raw image was processed only with debayering and white balancing using the grey card.
-(Figure 7 shows some examples).
-Figure 7: Some examples of pictures of the same table grape bunches were presented to agronomists on different occasions to test whether human experts consistently evaluated bunch ripeness, regardless of color correction, as described in Section 3.1.2
-We presented all 100 images in random sequence to a group of agronomist experts, without revealing that each bunch appeared twice.
-The experts labeled the color quality of each image using the color scale described in Section 3, without performing any other manipulation or consideration.
-The null hypothesis is that there is no difference in labeling standard and white-balanced images apart from some generic labeling noise.
-If we define the mean group labeling error as:
-$\mu_{i}=\mathbf{E}\left[L_{GT_{i}}-L_{exp_{i}}\right]$ (7)
-where $i\in[0,1]$ is the sample group, $L_{GT_{i}}$ is the hypothetical ground truth color label for a random sample from group $i$, and $L_{exp_{i}}$ is the label that an expert gives for the same random sample.
-Furthermore, we denote with $Y_{i}$ the sample mean of the difference between the true label and the estimated one.
-Then, we can state that $Y_{1}-Y_{0}$ is an unbiased normally distributed estimator of $\mu_{1}-\mu_{0}$, and thus we can formulate the following null hypothesis:
-$\mathcal{H}_{0}:\mu_{1}-\mu_{0}=0$ (8)
-which we can easily test for significance with a z-test.
-Analysis of the results revealed only a small difference between expert labels for the two groups, with a mean difference of $0.1$ and a standard deviation of $0.50$.
-The paired test for the sample batch gives a z-statistic of 1.4 (p=0.17), which does not provide sufficient evidence to reject $\mathcal{H}_{0}$.
-While a larger sample size might yield more definitive results, we chose instead to focus on the estimation algorithms, which are the main subject of our study.
-In particular, we included the comparison between white-balanced images and standard images into the cross-validation of the histogram-based features to gain deeper insights.
-3.2 Histogram based color and °Brix estimation
-This Section first outlines the experimental setup used to evaluate histogram features and subsequently presents the results.
-3.2.1 Description
-Given the exploratory experiments described in the previous Sections 3.1.1 and 3.1.2, we designed the histogram features as described in Section 2.2 and use them to train regression and classification models to predict SSC and color level, respectively.
-The following experiments were conducted for both color estimators (multi-class classification) and SSC estimators (°Brix regression).
-Given the higher difficulty of SSC estimation compared to exterior color evaluation, the following description focuses on SSC estimation, while the tables and figures summarize the results for both SSC and color estimation.
-The main question addressed in these experiments is the overall performance of the histogram-based algorithms and the influence of hyper-parameters.
-A number of grid hyper-parameter searches using Ridge Regression were conducted on a limited dataset, specifically the $T1_{STD}$ and $T1_{WB}$ splits, to mitigate distribution shifts due to different camera sensors and seasons during the initial analysis.
-These splits contain 112 samples each, divided into 89 train images and 23 validation images.
-The training images were augmented with horizontal flips, resulting in a total of 178 training images.
-As previously mentioned, both standard JPG and white-balanced images were evaluated to assess the influence of post-processing on the feature-based algorithms.
-To set a baseline for this approach, we considered the intrinsic variability training and test data.
-The $T1$ datasets have a mean of 17.19 °Brix, a standard deviation of 2.43 °Brix, and a mean absolute deviation (MAD) of 1.99 °Brix.
-Appendix A provides more details on the statistical properties of the datasets.
-The naive predictor, which always predicts the mean °Brix values of the training set, resulting in a mean absolute error (MAE) equal to the MAD, serves as the lower bound for estimator performance.
-Figure 8: a) Correlation between validation error and hyper-parameters for the top 10% performing runs.
-This plot highlights the direction (positive or negative) of the correlation.
-b) Test error for models with hyper-parameters identical to the best model shown in Table 3, but with varying $alpha$ values and color correction.
-The consistent lower error for models trained on standard images compared to color-corrected images gives a clear answer to the question raised in Section 3.1.2.
-Test MAE Val MAE Train MAE $alpha$ $n\_bin\_x$ $n\_bin\_y$ $cross$ $color\_space$ $feature\_len$ $log(alpha)\times feature\_length$
-1.459898 2.098344 1.583124 9.000000 16 8 none hsv 128 122.143041
-1.509351 2.129067 1.548176 9.000000 8 16 none hsv 128 122.143041
-1.522208 2.141015 1.614265 3.000000 8 8 none hsv 64 30.535760
-1.509816 2.144949 1.178387 9.000000 8 32 none hsv 256 244.286082
-1.507421 2.164503 1.619756 1.000000 16 8 thin lab 80 0.000000
-1.495739 2.164640 1.254232 3.000000 16 8 none hsv 128 61.071521
-1.517718 2.181812 1.431866 3.000000 16 8 fat hsv 96 45.803640
-1.523156 2.193494 1.719893 0.330000 8 8 thin lab 40 -19.259442
-1.431197 2.200048 1.189874 3.000000 8 16 none hsv 128 61.071521
-1.491955 2.327332 1.225366 1.000000 16 8 thin hsv 80 0.000000
-Table 3: The top 10 histogram-based models, ranked by validation mean absolute error (Val MAE), along with their corresponding hyper-parameters values.
-Some trends discussed in Section 3.2, such as the importance of alternative color spaces like HSV, are clearly evident.
-3.2.2 Discussion
-A grid search on the aforementioned hyper-parameter ranges, requiring training 3888 different models, was performed using the Weights & Biases experiment tracking framework [2] to compare and analyze the results.
-Further details of the hyper-parameter search are provided in Appendix B, while this section focuses on the key findings.
-The correlation coefficient of each feature with the validation error has been computed and is shown in Fig.
-8(a).
-In addition, Table 3 shows the 10 best runs for the $low\_res$ images.
-As expected, given the limited number of samples in the $T1$ datasets, the regularization parameters and those influencing feature vector length had the greatest impact on the final results.
-Furthermore, image resolution proved irrelevant, suggesting the safe use of lower-resolution images for computational efficiency in embedded applications.
-This is consistent with the feature construction process, where binning and averaging naturally remove the complexities of high-level details.
-Finally, and most notably, both Fig.8(a) and Table 3, indicate that using an alternative color space, such as HSV, helps in finding better coefficients for the Ridge Regression.
-Focusing solely on the 1296 lower-resolution runs, the best-performing models have validation MAE of $1.45$, with different combinations of hyper-parameters that have in common a low feature count relative to the regularization value.
-Table 3 summarizes some of these top-performing runs, all exhibiting MAE lower than the baseline brixmad of 1.99.
-The best run reaches a $\mathbf{27\%}$ improvement over the baseline, a notable result considering the limited training data.
-Figure 8(b) displays the validation errors for different $alpha$ values (on a logarithmic scale), for both color-corrected and standard images.
-The most interesting feature of the plot, which is a common trend in all similarly produced plots, is the consistently inferior performance of the models trained on color-corrected images compared to their standard counterparts with identical hyper-parameters.
-3.3 Multi-task Learning DNN
-The SSC and color distribution can be viewed as conditioned on the environmental and agronomic parameters of the field we are considering.
-While both the color and the SSC depend on the internal chemical state of the berries, this state itself is influenced by environmental factors such as growing conditions, training practices, and nutrient availability, The overall appearance of the grape bunches is also affected by these factors, which we assume are homogeneous across the vineyard.
-The architecture of the multi-task network, illustrated in Fig.
-5, reflects this reasoning.
-The encoder component captures common features in appearance variations that may correlate with the internal chemical state of the grapes.
-These shared features are then leveraged by two separate prediction heads to estimate the color class and SSC in parallel.
-As before, in the following we first give a description of the experiment and then discuss the results.
-3.3.1 Description
-The $T2$ dataset (Table 1) is used to evaluate the covariate shift due to the variability in treatments between consecutive seasons while keeping the sensor type fixed.
-Since phone images were captured in both raw and JPG formats, the $T2$ splits include standard ($T2_{STD}$) and white-balanced ($T2_{WB}$) versions.
-This allows for a direct comparison of performances between standard JPG and color-corrected images, under identical hyper-parameters and for the same grape bunch instances.
-In both cases, data from the 2021 season serves as the training and validation sets, while the data from the 2022 season are used as the test set.
-The $T2_{ALL}$ split consists only of color-corrected images since it includes reflex camera images that have been captured in raw format and subsequently color-corrected using the gray card.
-This split helps assess to some degree the effect of using two different camera devices to capture the same instances with the same perspective.
-The $T3$ splits combine images from all devices across both seasons, including both color-corrected and standard versions.
-Appendix C and Table 10 summarize the hyper-parameter ranges considered for the experimental evaluation.
-We performed $4912$ runs on a DGX server with four $A100$ GPUs.
-Each run was trained for $200$ epochs with hyper-band early stopping.
-For all the experiments we used a ResNet18 backbone [7], exploring both the effect of pretraining on ImageNet [4] and training from scratch.
-Hyper-parameter relevance was assessed using the BCH metric described in Section 2.3.1.
-While the complete details on the analysis of the importance of each hyper-parameter are described in Appendix C, in the following Section we discuss the most relevant findings.
-Figure 9: Multi-task network hyper-parameter correlation - These histograms shows the correlation between the validation error and the hyper-parameters computed on the 25% best performing runs.
-a) $T2$ dataset: it can be noted that the main influence comes from the classical parameters that impact any DNN training (learning rate, batch size, regularization parameters), though image size and color space remain relevant.
-More insights on this can be extracted from Table 4.
-b) $T3$ dataset: similar trends are observed as in (a), with classical DNN training parameters having the most significant impact.
-The increased importance of blur parameters suggests the potential benefits of data augmentation strategies in mitigating the effects of covariate shifts due to variations in camera devices.
-3.3.2 Discussion
-We begin our analysis with the $1002$ runs performed on the $T2$ splits.
-Since we sampled a wide range of models via random search, we concentrate on the best $25\%$ of the total runs to discuss the influence of hyper-parameters.
-Figure 9(a) presents the correlation between the hyper-parameters and the BCH metric, while Table4 lists the top $10$ performing runs.
-The multi-task network effectively learns good models for both SSC and color estimation.
-Apart from the typical learning parameters, the ones that show more influence on the performance results are image size (particularly width) and color space (rarely RGB among top performers).
-However, no single hyper-parameter shows high correlation values, indicating a complex relationship linking the parameters to the error.
-Another consideration is that $\alpha$ is typically ten times larger than $\beta$, giving a rough estimate of the relative weight of the corresponding losses for the final results.
-Moreover, the benefit of the additional augmentations appears less substantial than initially anticipated.
-It is important to note that the lowest MAE for SSC and the highest accuracy for color estimation are not achieved at the same epoch.
-This is unsurprising, as perfectly correlated feature learning for distinct tasks is unlikely, even with strong underlying correlations.
-Therefore, when using a single feature extractor, a trade-off must be made between optimizing for SSC or color estimation.
-If computational resources permit, employing two different feature extractors (i.e., different checkpoints) allows for independent optimization and potentially better performance on both tasks.
-Val BCH Val SSC MAE Val Acc $alpha$ $beta$ $im\_width$ $im\_height$ $momentum$ $learning\_rate$ $batch\_size$ $blur\_\sigma\_min$ $blur\_\sigma\_max$ $pretrained$ $one\_cycle$ $HSV$ $Lab$ $augment\_extra$
-4.076824 1.555511 0.673913 500 50 256 256 0.990000 0.000001 64 0.100000 2 False False False True True
-4.188802 1.832541 0.663043 1000 1000 256 64 0.950000 0.000001 32 0.500000 4 False True False True False
-4.265176 1.653810 0.554348 100 10 64 256 0.950000 0.000050 64 0.500000 2 True False True False False
-4.312880 1.732056 0.619565 500 100 256 64 0.800000 0.000050 128 0.100000 2 True True True False True
-4.346299 1.579317 0.630435 1000 100 256 128 0.800000 0.000050 128 0.010000 4 False True True False False
-4.412035 2.160111 0.532609 500 100 64 128 0.990000 0.000050 128 0.010000 2 False True True False False
-4.453714 1.795929 0.684783 1000 500 256 256 0.950000 0.000010 128 0.010000 2 False False False True True
-4.496901 1.891276 0.673913 100 50 256 128 0.900000 0.000050 128 0.500000 2 True True True False True
-4.536293 1.629490 0.521739 100 100 256 128 0.900000 0.000050 64 0.500000 2 False False False False False
-4.598450 1.620318 0.695652 10 1 256 256 0.990000 0.000050 128 0.100000 4 False True False True True
-Table 4: The best ten runs on $T2$ split.
-Similar trends emerge from the experiments on the $T3$ splits.
-Figure 9(b) shows the correlations of the hyper-parameters with the error, and Table 13 displays the best-performing runs.
-Again, the $\alpha$ to $\beta$ ratio is confirmed, along with the general preference for having higher image width values.
-The $T3$ test set, captured with a Realsense d435i, introduces a covariate shift due to its lower resolution and different white balance settings compared to the phone and reflex cameras used for training.
-This difference might explain the increased relevance of blur-related parameters in the augmentation.
-Following the analysis conducted for histogram features in Section 3.2, we compare the best $T3$ run with runs using the same hyper-parameters but trained on white-balanced images.
-Unlike the histogram feature case with a single regularization parameter, the multi-task model utilizes two ($\alpha$ and $\beta$).
-Given what was observed from the experiments, that the ratio of the $\alpha$ and $\beta$ values is more important than their absolute value, we use this as a parameter to compare the impact of using color-corrected images versus using the standard JPG ones.
-Similarly to the analysis presented in Section 3.2, we select the 150 best-performing runs and the ones with matching hyper-parameters (except for $\log(\alpha/\beta)$) and use them to extract a general trend shown in the scatter plot in Figure 10.
-The regression lines for both standard and white-balanced image sets exhibit nearly identical slopes ($-0.212$ and $-0.216$, respectively), and the difference between the lines is small, with largely overlapping confidence intervals.
-This suggests that the multi-task model effectively extracts relevant features regardless of color correction, with performance oscillations likely stemming from training data variability and other hyper-parameters.
-To conclude, the best-performing runs for SSC estimation achieve MAE values as low as 1.55 (1.57) for the model trained on the $T2_{STD}$ ($T2_{WB})$ split, and 1.11 (1.10) for the one trained on the $T3_{STD}$ ($T3_{WB})$ split.
-Considering that the MAD values for the $T2$ and $T3$ test splits are 1.71 and 1.36, respectively, the performance gain of the multi-task models is $9.3\%$ ($8.1\%$) for the $T2$ case and $18.4\%$ ($19.1\%$) for $T3$ case.
-These results confirm the negligible impact of color correction on the multi-task model performance, as observed in the regression analysis.
-The superior performance of the white-balanced model on the $T3$ dataset can be attributed to the automatic white balancing performed by the d435i camera used for collecting the $T3$ test set.
-Figure 10: Linear regression analysis of runs with hyper-parameters similar to the top 10 runs on the $T3_{STD}$ and $T3_{WB}$ datasets.
-The shaded area represents the $95\%$ confidence interval.
-While the models trained on color-corrected samples present slightly higher values on average, the smaller difference compared to Fig.
-8(b) suggests no practical advantage in using standard images over the white-balanced ones in this context.
-4 Conclusions
-This work addressed the problem of estimating soluble solid content (SSC) in grapes using RGB data acquired in the field.
-While numerous studies explore laboratory equipment and hyperspectral data for SSC estimation through visual sensors, this study is among the few to focus exclusively on the practical scenario of RGB data collected directly in field conditions.
-The complexity of this task arises from covariate shifts inherent to the problem, including seasonal variations and differences in imaging devices.
-To tackle this challenge, a dataset spanning multiple years was collected, incorporating data from various camera devices and annotated for SSC and color quality based on agronomic practices.
-Human performance in assessing the correlation between color and SSC was evaluated through two user studies, which informed the design of data-driven algorithms for estimating both SSC and color.
-Two algorithmic approaches were explored: a light-weight method using classical handcrafted features, suitable for devices with limited computational power, and a multi-task neural network architecture that, while more computationally demanding, effectively leverages correlations between visual appearance (color and shades) and SSC.
-Both methods were experimentally validated using field-collected data from the Pizzutello Nero grape variety.
-Results indicate that the proposed techniques achieve reliable SSC estimation with a reasonable margin compared to the average human estimator.
-Furthermore, experiments demonstrated some robustness to covariate shifts, including seasonal variations and device differences, positioning the proposed solutions as viable prototypes for low-cost, on-field applications where frequent data collection and labeling are challenging.
-We release all the data, code, trained models, and training scripts used in this study.
-Future work will focus on refining and aligning these low-cost estimators and benchmarking them against hyperspectral approaches.
-Acknowledgments
-[…]
-This work is part of a project that has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No.
-101016906 – Project CANOPIES
-[…]
-This work has been partially supported by project AGRITECH Spoke 9 - Codice progetto MUR: AGRITECH ”National Research Centre for Agricultural Technologies” - CUP CN00000022, of the National Recovery and Resilience Plan (PNRR) financed by the European Union ”Next Generation EU”.
-This work has been partially supported by Sapienza University of Rome as part of the work for project H&M: Hyperspectral and Multispectral Fruit Sugar Content Estimation for Robot Harvesting Operations in Difficult Environments, Del.
-SA n.36/2022.
-Appendix A Detailed dataset description
-Data collection was part of a larger effort within the Canopies project, exploring different sensors across seasons.
-The first data collection was done in September-October 2021, using a smartphone camera and a Reflex camera.
-The second data collection took place in September 2022, utilizing a smartphone camera and a Realsense d435i.
-Although the d435i is an RGB-D device, we used it as a standard camera for °Brix and color estimation.
-Table 5 summarizes the different types of data collected during the two seasons.
-The naming of the splits in the following tables follows a descriptive naming used during the development of the dataset and the code.
-For ease of reference, we renamed the splits with simpler tags ($T1$, $T2$, $T3$) to avoid bloated experiment descriptions.
-The match between the splits names given in Table 1 and 7 are the following:
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-The data organization allows for testing covariate shifts due to both sensor type and season.
-These atomic splits represent minimal groups sharing the same sensor, season, image type, and label characteristics.
-Using these atomic splits, we designed the experimental splits, summarized in Table 6 and Table 7.
-Images: Smartphone images were acquired at a resolution of 3000x4000 pixels, in both JPG and raw formats, with a few exceptions.
-Reflex images, captured in raw format, have a resolution of 3900x2613 pixels.
-The d435i RGB images have a lower resolution of 1280x720 pixels.
-Raw images were corrected for white balance using a gray card (Fig.2).
-Not all images included a gray card due to collection errors, but these images were still used for comparison purposes.
-The full images were labeled with bounding boxes around the grape bunches and then cropped.
-The resulting cropped images have very different sizes since some images were close-ups while others contained multiple targets.
-The combination of season, sensor used, gray card presence, and other factors creates differences summarized in Table 5.
-The rows in the table represent atomic splits, i.e., splits with the same distribution.
-In contrast, between the splits, there is some form of covariate shift that we aim to study.
-Datset Id Year Smartphone Reflex d435i Num images Num samples Expert GT Color GT
-Type Jpg Raw Card Raw Card RGB IR Card
-A 2021 motog8 yes no no no - no no - 100 100 no no
-B 2021 motog8 no yes yes yes yes no no - 38 38 possible yes
-C 2021 motog8 yes yes yes yes yes no no - 62 62 possible yes
-D 2021 motog8 yes yes yes yes yes no no - 50 50 possible yes
-E 2022 realmi yes yes no no - yes no no 1 2 no no
-F 2022 realmi yes yes yes no - yes no yes 6 22 possible yes
-G 2022 realmi yes yes no no - yes no yes 1 2 no no
-H 2022 realmi yes yes yes no - yes no yes 4 19 possible yes
-I 2022 motog8 yes yes yes no - yes yes yes 8 51 possible yes
-Table 5: Description of the atomic splits according to the device, image format, gray card presence, and the possibility of having expert ground truth from white-balanced images.
-Name Components Description
-2021PhoneJPG A+C+D Contains all the jpg images collected with a smartphone in 2021, without white balance from the spits A, C, and D.
-2021PhoneJPG_reduced C+D Contains the jpg images collected with a smartphone in 2021 that have a corresponding white-balanced image. Corresponds to splits C and D.
-2021PhoneJPG_WB_reduced C+D Contain all the white-balanced images collected with a smartphone in 2021. Corresponds to splits C and D.
-2021PhoneJPG_WB B+C+D Contains all the jpg images collected with a smartphone in 2021, from splits B, C, and D
-2021ReflexJPG_WB B+C+D Contains all the jpg images collected with a reflex camera in 2021 from splits B, C, and D
-2022PhoneJPG_WB F+H+I Contains all the jpg images collected with a smartphone in 2022 from splits F, H, and I
-2022d435iAutoWB F+H+I Contains all the auto white balanced jpg images collected with d435i in 2022, from splits F, H, and I
-Table 6: Description of the basic splits used to train and test the proposed algorithms.
-Name Train set components Test set components Description
-Season21ToSeason22Phone A+C+D F+H+I All jpg images taken with a phone.
-Season21ToSeason22Phone_WB B+C+D F+H+I All raw images taken with a phone that have been white-balanced.
-Season21ToSeason22All_WB B+C+D F+H+I All phone and reflex images white-balanced.
-Phone_d435i_Std A+C+D+F+H+I All phone images not white-balanced, plus all d435i images
-AllWB_d435i_Std B+C+D+F+H+I All phone and reflex images white-balanced, plus all d435i images
-MixedSeason_JPG A+C+D+H+I F All jpg images, keeping for testing only the ones with the most uncertain brix value (around 16).
-MixedSeason_WB_WB B+C+D+H+I F All white balanced images, keeping for test only the most uncertain brix value (around 16).
-Table 7: This table describes the most complex data splits, which we called mixed splits. These splits span across different sensors, while the train and test separation is based on the season.
-Split name #$\#samples$ $Brix_{mean}$ $Brix_{std}$ $Brix_{mae}$ $CC_{1}$ $CC_{2}$ $CC_{3}$ $CC_{4}$ $CC_{5}$ $CC_{6}$
-2021PhoneJPG 212 16.23 2.52 2.02 0.01 0.04 0.09 0.14 0.27 0.45
-2021PhoneJPG_WB 150 17.54 2.46 2.04 0.01 0.03 0.06 0.10 0.21 0.59
-2021PhoneJPG_reduced 112 17.19 2.43 1.99 0.01 0.04 0.07 0.12 0.22 0.54
-2021PhoneJPG_WB_reduced 112 17.19 2.43 1.99 0.01 0.04 0.07 0.12 0.22 0.54
-2022PhoneJPG 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
-2022PhoneJPG_WB 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
-2021ReflexJPG_WB 150 17.54 2.46 2.04 0.01 0.03 0.06 0.10 0.21 0.59
-2022d435iAutoWB 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
-MixedSeason_train 352 17.67 2.75 2.33 0.01 0.04 0.12 0.16 0.32 0.34
-MixedSeason_test 44 15.96 1.68 1.36 0.00 0.00 0.09 0.09 0.27 0.55
-MixedSeason_WB_train 440 18.27 2.40 1.98 0.01 0.04 0.10 0.13 0.27 0.46
-MixedSeason_WB_test 44 15.96 1.68 1.36 0.00 0.00 0.09 0.09 0.27 0.55
-Season21ToSeason22Phone_train 212 16.23 2.52 2.02 0.01 0.04 0.09 0.14 0.27 0.45
-Season21ToSeason22Phone_test 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
-Season21ToSeason22Phone_WB_train 150 17.54 2.46 2.04 0.01 0.03 0.06 0.10 0.21 0.59
-Season21ToSeason22Phone_WB_test 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
-Season21ToSeason22All_WB_train 300 17.54 2.46 2.04 0.01 0.03 0.06 0.10 0.21 0.59
-Season21ToSeason22All_WB_test 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
-All_Std 396 17.48 2.71 2.30 0.01 0.04 0.12 0.15 0.32 0.37
-All_Std_WB 484 18.06 2.44 2.03 0.01 0.03 0.10 0.12 0.27 0.47
-Table 8: Basic statistics for all the atomic and mixed splits described in Table 7.
-For each split the following statistics are listed: number of samples, average °Brix, standard deviation of the °Brix, °Brix mean absolute error, and the percentage of each color class in the dataset.
-The color classes are unbalanced towards the unripe color classes, since the ripe bunches were harvested every few days.
-Appendix B Hyper-parameters Exploration for the Histogram based Estimators
-Following from Section 3.2, here we give the details of the hyper-parameter search for the histogram features.
-Table 9 summarizes the hyper-parameter ranges investigated by this experiment.
-Most of them have been introduced in Section 2.2.
-In addition to those, here we add the investigation of the effect of the input image resolution (low_res: $320\times 240$, med_res: $640\times 480$, high_res: $1280\times 720$).
-For the $cross$ parameter, to avoid hyper-parameter explosion, we designed three possible configurations (none, thin, fat), as described in Section 2.2.
-Hyper-parameter Range description
-$n_{bin_{x}}$ 4, 8, 16 number of image columns bins.
-$n_{bin_{y}}$ 8, 16, 32 number of image rows bins.
-$cross$ ”none”, ”thin” or ”fat” How many bins to drop for background elimination. See Eq.1 for details.
-$color\_space$ ”none”, ”hsv” or ”Lab” Use of CIE-La*b*, HSV or RGB color coordinates.
-$dataset$ low_res, med_res, high_res, Image resolution, corresponding to $320\times 240$, $640\times 480$ and $1280\times 720$ respectively, in color corrected and standard form.
-wb_low_res, wb_med_res, wb_high_res
-$\alpha$ 0.33, 1, 3, 9, 27, 81, 243, 729 Ridge Regression regularization parameter.
-Table 9: Hyper-parameters ranges for the histogram feature based models.
-To identify the most relevant hyper-parameters, taking into account the non-linear interactions between the different features, we opted to train a random forest (RF) algorithm on the hyper-parameter values themselves, to predict the validation error.
-Using this technique allows for reading the importance score of the RF as an indicator of the relevance of each hyper-parameter [5].
-Figure 11(a) shows the importance factors for the RF regressor trained to predict the validation error from the hyper-parameter values.
-The importance factor show which features contributes the most to the RF regression of the validation error, allowing for concluding that the different starting image resolutions have little influence on the SSC estimation.
-This is a relevant finding, since it allow us to safely work with the lowest resolution images without trading for estimation accuracy.
-Among the most important variables we find the ones that control the model flexibility, i.e.
-the regularization parameter, the $x$ and $y$ bin number, and the cross parameters, that impact the feature vector length.
-Together these parameters have to balance prediction power with the risk of over-fitting.
-Finally, the color space show some relevance too.
-The correlation coefficient has already been shown in Fig.
-8(a), and its replicated in Fig.
-11(b) for convenience.
-Image resolution was omitted and the total histogram length ($feature\_len$) was added.
-That histogram helps understanding which variable linearly correlates with the validation error, in positive or negative way, but alone it’s not enough to show feature importance, since there could be some interactions effects between them.
-Figure 11: a) Training a Random Forest using the hyper-parameters values to predict the validation error and then extracting the importance scores is a common way to extract the relative impact of each feature on the final estimator performance.
-Here we can see that the choice of initial image resolution is not relevant, while the regularization, the number of bins (that influence feature length and flexibility), and the color space have to carefully selected.
-b) shows the correlation computed on the 10% best performing runs.
-This plot complements the other, showing where the correlation with the error is positive or negative.
-Figure 12: To better extract the impact of feature length on validation error, we add three parameters to the analysis, $feature_{l}en$ and its interactions with $alpha$ and $\log{alpha}$.
-It is easy to see that the length of the histogram is the most important hyper-parameter to be tuned for final model performance.
-To further explore this influence of feature flexibility due to histogram length, we show in Figure 12 the correlation between the validation error and the most important hyper-parameters extracted from Figure 11(a), explicitly adding $feature\_lenght$ and its interactions with $alpha$ and $\log(alpha)$.
-It is evident that the feature length is the most important factor among the others, and that the interaction between feature length and alpha is more than relevant.
-The general behavior that we can extract from this analysis is that the more are the bins of the histogram, the higher should be the alpha value, and together these two parameters are sufficient to explain a great deal of the validation error variability.
-Appendix C Hyper-parameters Exploration for the Multi-task DNN
-This appendix follow the same principles described in the previous one.
-Table 10 gives a summary of the hyper-parameters ranges explored in the sweep.
-Given the huge parameter space, this time the sweep was performed with random sampling, with a total run count of 4912 (see Section 3.3).
-Figures 13(a) and 14(a) show importance values for the RF trained on the hyper-parameter values to predict validation BCH respectively for $T2_{STD}$ (Season21ToSeason22_Phone) and $T3_{STD}$ (MixedSeason).
-Finally, Tables 4, 11, 13, 14, and 12 summarize the best 10 runs respectively on $T2_{STD}$, $T2_{WB}$, $T3_{STD}$, $T3_{WB}$ and $T2_{ALL}$.
-Multi-task learning hyper-parameter Range description
-$\alpha$ and $\beta$ 1, 10, 50, 100, 500, 1000 The loss balancing factors (Eq. 2)
-$batch\_size$ 32, 64, 128 Training batch size
-$learning\_rete$ 0.00001, 0.00005, 0.000001 Initial learning rate, to be modified by the scheduling strategy
-$learning\_rete$ 0.00001, 0.00005, 0.000001 Initial learning rate, to be modified by the scheduling strategy
-$momentum$ 0.8, 0.9, 0.95, 0.99 Momentum for the SGD optimizer
-$im\_height$ and $im\_width$ 64, 128, 256 The image starting sizes before feature extraction by the auto-encoder.
-$blur\_\sigma_{min}$ and $blur\_\sigma_{max}$ $\left[0.01,0.1,0.5\right]$ and $\left[2.0,4.0\right]$ Together they define the range for the Gaussian blur augmentation that is always applied.
-$color\_space$ ”none”, ”hsv” or ”Lab” Use of CIE-La*b*, HSV or RGB color coordinates.
-$one\_cycle$ $True$ or $False$ Whether to use of not the one cycle scheduling strategy
-$pretrained$ $True$ or $False$ Whether the resnet18 backbone is pretrained on ImageNet or not.
-$dataset\_split$ $Season21ToSeason22Phone$, $Season21ToSeason22Phone\_WB$, $Season21ToSeason22All\_WB$, $MixedSeason$, $MixedSeason\_WB$ Image resolution, corresponding to $320\times 240$, $640\times 480$ and $1280\times 720$ respectively, in color corrected and standard form.
-$\alpha$ 0.33, 1, 3, 9, 27, 81, 243, 729 Ridge Regression regularization parameter.
-Table 10: This Table summarizes the hyper-parameters descriptions for the Multi-task experiments.
-A wide range of hyper-parameters was explored using random search for a total of more than $5000$ experimental runs.
-In this case, the dataset splits used are not to be considered a proper hyper-parameter, but as a defining subset of the parameter sweep, since the distribution shifts between them are considerable.
-They are included in this table for ease of reference.
-Figure 13: a) These plots show the importance and correlation factor of the best $25\%$ of the runs trained on the $Season2021ToSeason2022$ dataset split.
-It can be noted that the main influence comes from the classical parameters that impact any DNN training (learning rate, batch size, regularization parameters), while the choice of image size and color space comes seems to be still relevant for the final performances.
-More insights on this can be extracted from Table 4.
-Figure 14: These plots show the importance and correlation factor of the best $25\%$ of the runs trained on the $MixedSeason$ dataset split.
-Val BCH Val SSC MAE Val Acc $alpha$ $beta$ $im\_width$ $im\_height$ $momentum$ $learning\_rate$ $batch\_size$ $blur\_\sigma\_min$ $blur\_\sigma\_max$ $pretrained$ $one\_cycle$ $HSV$ $Lab$ $augment\_extra$
-4.100699 1.835277 0.478261 500 50 128 256 0.800000 0.000050 64 0.500000 2 True True True False True
-4.132786 1.913489 0.576087 100 1 256 64 0.990000 0.000050 64 0.010000 2 True True False False True
-4.384027 1.570156 0.554348 1000 10 256 256 0.990000 0.000001 32 0.010000 2 True False False False True
-4.397698 1.605315 0.434783 500 10 128 128 0.800000 0.000050 32 0.500000 2 False True False True True
-4.499346 1.982949 0.576087 100 100 256 256 0.950000 0.000050 64 0.010000 4 True True True False False
-4.522784 1.714363 0.500000 1000 100 256 128 0.990000 0.000010 64 0.100000 2 True False False True False
-4.552584 1.768912 0.489130 500 10 128 128 0.900000 0.000050 64 0.500000 2 True True False True False
-4.558001 1.644159 0.445652 1000 500 256 128 0.800000 0.000010 32 0.100000 4 False True False True False
-4.583944 1.916520 0.521739 500 10 256 256 0.950000 0.000010 64 0.500000 4 False False False False True
-4.596296 1.968404 0.597826 100 1 64 64 0.990000 0.000050 32 0.100000 4 True True True False True
-Table 11: The best ten runs on $Season2021ToSeason2022\_WB$ split. Compared to the ones given in Table 4, both the SSC MAE estimation error and Color estimation accuracy show worse performance.
-Val BCH Val SSC MAE Val Acc $alpha$ $beta$ $im\_width$ $im\_height$ $momentum$ $learning\_rate$ $batch\_size$ $blur\_\sigma\_min$ $blur\_\sigma\_max$ $pretrained$ $one\_cycle$ $HSV$ $Lab$ $augment\_extra$
-3.977163 1.754222 0.576087 50 1 128 128 0.990000 0.000050 32 0.500000 2 True True False True True
-4.123796 1.945782 0.532609 100 10 128 128 0.950000 0.000050 64 0.010000 2 False False True False False
-4.269722 1.610763 0.673913 500 500 256 256 0.990000 0.000010 64 0.100000 4 True True True False True
-4.307943 1.754948 0.608696 500 10 128 256 0.950000 0.000010 32 0.100000 2 True True False False False
-4.327728 1.697267 0.467391 1000 50 128 128 0.950000 0.000010 32 0.500000 4 False True False True False
-4.336076 1.930476 0.532609 1000 10 64 256 0.800000 0.000050 64 0.010000 2 True False False True False
-4.347323 1.688515 0.500000 500 100 256 256 0.900000 0.000010 64 0.100000 4 False False False True True
-4.411548 1.531855 0.597826 500 50 256 256 0.990000 0.000010 32 0.500000 2 False False False False False
-4.440626 1.685545 0.630435 1000 500 256 256 0.990000 0.000001 128 0.100000 2 False False True False False
-4.501309 1.803873 0.608696 1000 10 256 256 0.800000 0.000050 64 0.100000 4 True False False True True
-Table 12: The best ten runs on $Season2021ToSeason2022All\_WB$ split.
-Val BCH Val SSC MAE Val Acc $alpha$ $beta$ $im\_width$ $im\_height$ $momentum$ $learning\_rate$ $batch\_size$ $blur\_\sigma\_min$ $blur\_\sigma\_max$ $pretrained$ $one\_cycle$ $HSV$ $Lab$ $augment\_extra$
-1.581554 1.200569 0.795455 500 10 128 64 0.900000 0.000050 32 0.010000 2 True False True False True
-1.662561 1.116875 0.772727 1000 50 256 128 0.990000 0.000010 32 0.100000 4 True False False False True
-1.686950 1.162343 0.772727 1000 50 256 256 0.950000 0.000010 32 0.010000 4 True True False True False
-1.723328 1.231537 0.795455 50 100 256 256 0.950000 0.000010 32 0.010000 4 False False True False True
-1.741897 1.102953 0.795455 1000 1000 64 64 0.990000 0.000001 32 0.500000 4 False True False True True
-1.745264 1.301862 0.795455 1000 1000 128 128 0.950000 0.000001 32 0.100000 4 False True False False False
-1.746910 1.269346 0.818182 1000 1000 256 256 0.900000 0.000010 128 0.100000 2 True True False True False
-1.771968 1.174211 0.795455 100 500 64 128 0.950000 0.000010 64 0.100000 4 False False False False True
-1.788137 1.190380 0.772727 100 50 128 256 0.950000 0.000050 128 0.500000 4 False False False False False
-1.788171 1.242672 0.818182 10 10 256 256 0.950000 0.000050 64 0.100000 4 True True False False True
-Table 13: The best ten runs on $MixedSeason$ ($T3_{STD}$) split.
-The validation set is the same as for the $MixedSeason\_WB$ ($T3_{WB}$), that is collected with a d435i camera.
-Since the images from this device are automatically white balanced, this imply that the model faces a strong covariate shift due to different camera device and image modality.
-Val BCH Val SSC MAE Val Acc $alpha$ $beta$ $im\_width$ $im\_height$ $momentum$ $learning\_rate$ $batch\_size$ $blur\_\sigma\_min$ $blur\_\sigma\_max$ $pretrained$ $one\_cycle$ $HSV$ $Lab$ $augment\_extra$
-1.545861 1.307749 0.818182 50 100 64 128 0.800000 0.000050 32 0.500000 4 True True False False False
-1.596429 1.101236 0.795455 100 100 128 256 0.800000 0.000010 128 0.010000 2 False False True False False
-1.611784 1.219757 0.772727 500 1000 128 128 0.800000 0.000010 32 0.010000 4 False False True False False
-1.639119 1.065401 0.727273 50 10 128 256 0.950000 0.000050 64 0.500000 2 False True False False True
-1.688067 1.220433 0.750000 500 500 128 128 0.950000 0.000010 32 0.500000 4 True False True False False
-1.709684 1.295742 0.704545 10 10 256 256 0.990000 0.000010 32 0.010000 2 True False False True False
-1.715017 1.159092 0.772727 100 100 64 128 0.800000 0.000050 32 0.100000 2 True False True False False
-1.715871 1.158179 0.840909 10 10 256 256 0.900000 0.000050 64 0.100000 2 True False True False True
-1.759702 1.200407 0.795455 50 50 128 128 0.900000 0.000050 128 0.100000 4 False False True False False
-1.776299 1.200725 0.772727 100 50 256 256 0.900000 0.000050 32 0.010000 4 True True True False False
-Table 14: The best ten runs on $MixedSeason\_WB$ ($T3_{WB}$)split.
-This time the results are comparable with the STD ones presented in Table 13.
-This can be explained considering the fact that the validation images of the d435i are white-balanced and therefore with a similar hue to the color-corrected training images.